### v1

# A quasilinear Keller-Segel model with saturated discontinuous advection

Abstract.
We consider the singular limit of a chemotaxis model of bacterial collective motion recently introduced in [2]. The equation models aggregation-diffusion phenomena with advection that is discontinuous and depends sharply on the gradient of the density itself. The quasi-linearity of the problem poses major challenges in the construction of the solution and complications arise in the proof of regularity. Our method overcomes these obstacle by relying solely on entropy inequalities and the theory of monotone operators. We provide existence, uniqueness and smoothing estimates in any dimensional space.
M.P.G. is partially supported by NSF Grant DMS-2206677. The authors would like to thank Franca Hoffmann and Inwon Kim for the fruitful discussions.
1. Introduction
Since the classical experiment of J. Adler on Escherichia coli [1] many experiments have confirmed the formation and propagation of traveling pulses of bacteria. The Keller-Segel model has been widely used as a basis to explain this phenomena in bacterial colonies driven by chemotaxis, see [13, 14]: it describes how the cell density moves towards a nutrient $N$ and a chemical molecule $S$ (chemoatractant) which is generated by the cells themselves. The velocity of this movement is proportional to the gradient of these two quantities. The Keller-Segel model has been the subject of intense mathematical investigation.
In recent works ([3, 15, 11, 7, 4, 6]), a novel class of models has been formulated, originating from a mesoscopic description of the run-and-tumble phenomenon at the individual cell level. Mathematically, these models deviate from the classical Keller-Segel framework in that the cells’ velocity (or flux) is saturated and is proportional to $\frac{\nabla S}{|\nabla S|}$ and $\frac{\nabla N}{|\nabla N|}$. The model, in its more general form, reads as
$\partial_{t}\rho$ $=D_{\rho}\Delta\rho-\textrm{div}(\rho u_{S}+\rho u_{N}),$
$\partial_{t}S$ $=D_{S}\Delta S-\alpha S+\beta\rho,$
$\partial_{t}N$ $=D_{N}\Delta N-\gamma\rho N,$
where
$u_{S}$ $=\chi_{S}J(\partial_{t}S,|\nabla S|)\frac{\nabla S}{|\nabla S|},\quad u_{N}=\chi_{N}J(\partial_{t}N,|\nabla N|)\frac{\nabla N}{|\nabla N|}.$
In [3], the authors present an explicit formula for traveling pulses and establish the existence of a steady state, mirroring the physical occurrence of cluster formation at the bacteria level (e.g., in the absence of nutrients). They also demonstrate the alignment of numerical experiments with actual data. The bounded fluxes, distinct from the flux in Keller-Segel, offer a significant advantage in investigating global well-posedness and, notably, in exploring traveling pulses and long-term asymptotics. Overall, this model has demonstrated greater robustness in replicating the observed traveling bands in experiments.
More recently, Calvez and Hoffmann [2] considered a simplified version of the above model (absence of nutrient and quasi-stationary chemoattractant signal)
$\left\{\begin{array}[]{rcll}\partial_{t}\rho-\Delta\rho+\chi\textrm{div}\,\left(\rho\frac{\nabla S}{|\nabla S|}\right)&=&0,\\ \vspace{2mm}-\Delta S+\alpha S&=&\rho,\end{array}\right.$ (3)
and perform nonlinear stability analysis. In particular they show that if the initial data is near, in some norm, to the equilibrium function $\rho_{\infty}=\frac{\chi}{2}e^{-\chi|x-x_{0}|}$, then the corresponding solution converges to $\rho_{\infty}$ exponentially fast.
In this manuscript we deal with the model obtained from (3) by performing the singular limit in $\alpha$. After rescaling $S\mapsto S/\alpha$, the limit $\alpha\rightarrow\infty$ yields the following quasi-linear parabolic problem:
$\partial_{t}\rho-\Delta\rho+\chi\textrm{div}\,\left(\rho\frac{\nabla\rho}{|\nabla\rho|}\right)=0.$ (4)
Unlike (3), the advection depends on the gradient of the density, making it a competing term with the diffusion. Note in fact that
$\Delta\rho-\chi\textrm{div}\,\left(\rho\frac{\nabla\rho}{|\nabla\rho|}\right)=\textrm{div}\,\left(\nabla\rho\left(1-\frac{\chi\rho}{|\nabla\rho|}\right)\right),$
which shows a negative diffusion for profiles that have $\chi\rho\geq|\nabla\rho|$. For example, Gaussian functions $e^{-|x|^{2}}$ will make the diffusion coefficient negative around the zero. The system is diffusive as long as $\lvert\nabla\rho\rvert>\chi\rho$, and the steady-state is governed by the equation
$\lvert\nabla\rho_{\infty}\rvert=\chi\rho_{\infty}.$ (5)
One solution to (5) is
$\rho_{\infty}(x)=ce^{-\chi|x-x_{0}|},$ (6)
for some $c_{1}$, $x_{0}$ determined by the conservation of mass and property of invariance by translation of the equation. Other admissible (i.e. belonging to $L^{1}(\mathbb{R}^{d})$) solutions to (5) exist; for example, for $d=1$ another solution is the “multi-peak profile”
$\rho_{\infty}(x)=\max\{c_{0}\exp(-\chi|x-x_{0}|),c_{1}\exp(-\chi|x-x_{1}|),...c_{k}\exp(-\chi|x-x_{k}|)\},$ (7)
for a suitable choice of $k\geq 0$, $c_{0}\ldots c_{k}>0$, and $x_{0}\ldots x_{k+1}\in\mathbb{R}$, while for $d\geq 2$ another solution is the “factorized profile”
$\rho_{\infty}(x)=c\exp\left(-\frac{\chi}{\sqrt{d}}\sum_{i=1}^{d}|x_{i}-x_{0,i}|\right),$ (8)
which equals the product of (suitably rescaled) single variable steady states. The proof of existence of solution to (4) is still out of reach. The main difficulty is the control on the diffusion coefficient, which can change along the flow and destroy the parabolic nature of the equation. We believe that if the initial data makes the diffusion coefficient positive, the coefficient will remain non-negative throughout the evolution. This is, however, still a claim. Problem (4) has a more tractable form in one dimension
$\partial_{t}\rho-\partial_{xx}\rho+\chi\partial_{x}\left(\rho\;\textrm{sign}(\rho_{x})\right)=0,$
and is the subject of current investigation [10].
As a first attempt to understand (4), we introduce an alternative model that retains some of the essential characteristics of (4) while being mathematically more tractable. Specifically, we consider
$\left\{\begin{array}[]{rcll}\partial_{t}\rho-\textrm{div}\,\left[\left(1-\frac{\chi\rho}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho\right]&=&0,&\quad\mbox{in }\mathbb{R}^{d}\times(0,\infty),\\ \vspace{2mm}\rho(0)&=&\rho_{in},&\quad\mbox{in }\mathbb{R}^{d},\end{array}\right.$ (9)
where $(\cdot)_{+}$ denotes the positive part. Here we use the convention $f=f_{+}-f_{-}$. Equation (9) is quasilinear. The type of nonlinearity poses major challenges in the construction of solutions. Let us, however, quickly summarize why global well-posedness for (9) is more in-reach. First, observe that the operator
$\left(1-\frac{\chi}{\lvert\nabla\log\rho\rvert}\right)_{+}\nabla\log\rho$
is monotone. More precisely,
$\left[\left(1-\frac{\chi}{\lvert\nabla\log\rho\rvert}\right)_{+}\nabla\log\rho-\left(1-\frac{\chi}{\lvert\nabla\log z\rvert}\right)_{+}\nabla\log z\right]\left(\nabla\log\rho-\nabla\log z\right)\leq 0.$
This monotonicity property, coupled with the theory of monotone operators and a fixed-point argument, proves fundamental in the proof of existence of solutions to (9) equipped with an extra viscosity term $\varepsilon\Delta\rho$. The crux of the analysis is the vanishing viscosity limit $\varepsilon\to 0$ and the corresponding identification of the limit of the nonlinear term
$\left(1-\frac{\chi}{\lvert\nabla\log\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}.$
A common strategy would involve showing strong convergence of one of the two terms. However, standard regularity techniques, usually based on obtaining estimates for the first derivative through second-order test functions, prove unsuitable here due to the discontinuity of the advection term and the discontinuity of $\nabla\rho$ itself. We overcome this problem using the Boltzmann relative entropy between two viscosity solutions. Let $H[\rho_{\varepsilon},\rho_{\delta}]$ be the nonnegative functional defined as
[…]
This functional measures the distance, in $L^{1}$ norm, between the functions $\rho_{\varepsilon}$ and $\rho_{\delta}$. Direct computations show that
$\frac{d}{dt}H[\rho_{\varepsilon},\rho_{\delta}]=$ $-\frac{1}{2}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}-\left(1-\frac{\chi\rho_{\delta}}{|\nabla\rho_{\delta}|}\right)_{+}\right]^{2}\,{\rm d}x$
$-\frac{1}{2}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left|\nabla\log\left(\frac{\rho_{\varepsilon}}{\rho_{\delta}}\right)\right|^{2}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}+\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\,{\rm d}x$
[…]
where the error terms derive from the viscosities $\varepsilon\Delta\rho_{\varepsilon}$ and $\delta\Delta\rho_{\delta}$. As $\varepsilon$ and $\delta$ approach zero, the error terms vanish, and consequently
$-\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}-\left(1-\frac{\chi\rho_{\delta}}{|\nabla\rho_{\delta}|}\right)_{+}\right]^{2}\,{\rm d}x\,{\rm d}t\to 0,$
and
$-\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left|\nabla\log\left(\frac{\rho_{\varepsilon}}{\rho_{\delta}}\right)\right|^{2}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}+\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\,{\rm d}x\,{\rm d}t\to 0.$
Note that $H[\rho_{\varepsilon},\rho_{\delta}](0)=0$ because $\rho_{\varepsilon}$and $\rho_{\delta}$ have same initial data. The two limits above will imply that, as $\varepsilon$ and $\delta$ approach zero,
$\left(1-\frac{\chi}{\lvert\nabla\log\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\rightharpoonup\left(1-\frac{\chi}{\lvert\nabla\log\rho\rvert}\right)_{+}\nabla\rho,$
weakly in $L^{2}$. The identification of this limit is the content of Section 4.1. Almost automatically, the relative entropy argument will also provide uniqueness of the solution for the system without viscosity. In conclusion, our analysis establishes the existence of a unique solution in $L^{p}(t,T,C^{\alpha}(\mathbb{R}^{d}))$ for any $t>0$. We obtain regularity for $\partial_{t}\rho$ in Sobolev spaces with negative coefficients. The question of whether one can obtain more regularity is left for future investigation. Potential insights might be drawn from existing results on the $p$-Laplacian equation [5, 8]. However, the structure of our degenerate set is slightly more complicated than the one of the $p$-Laplacian, namely $|\nabla\rho|=0$. A De Giorgi argument for Hölder regularity might pose challenges, since our diffusion coefficient is degenerate, probably in a set of positive measure, and depends in a nonlinear way on $\nabla\log\rho$.
Here is our main result:
Theorem 1.
Let $\rho_{{\rm in}}\in L^{p}(\mathbb{R}^{d})\cap L^{1}_{2}(\mathbb{R}^{d})$, with $p>d$, $\rho_{{\rm in}}>0$ a.e. in $\mathbb{R}^{d}$ and such that $\int_{\mathbb{R}^{d}}\rho_{{\rm in}}|\log{\rho_{{\rm in}}}|<\infty$. Then, for every $T>0$ there exists an unique solution $\rho:[0,T)\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ to (9) such that
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\partial_{t}\rho\,\varphi\,\,{\rm d}x\,{\rm d}t=-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho\nabla\varphi\,\,{\rm d}x\,{\rm d}t\quad\mbox{for }\;\;\forall\varphi\in L^{2}(0,T;H^{1}(\mathbb{R}^{d})),$ (10)
with $\rho\to\rho_{{\rm in}}$ strongly in $L^{p}(\mathbb{R}^{d})$ as $t\to 0$. This solution belongs to $C(0,T,L^{p}(\mathbb{R}^{d})\cap L^{p}(t,T,C^{\alpha}(\mathbb{R}^{d}))$ for all $t>0$ and has the following properties:
  1. (i)
    Mass Conservation:
    $\|\rho(t)\|_{L^{1}(\mathbb{R}^{d})}=\|\rho_{{\rm in}}\|_{L^{1}(\mathbb{R}^{d})}.$
  2. (ii)
    $L^{p}$ norm bound:
    $\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}\leq\|\rho_{{\rm in}}\|_{L^{p}(\mathbb{R}^{d})}.$
  3. (iii)
    Moment evolution:
    $\int_{\mathbb{R}^{d}}\rho(t)(1+|x|^{2})\,{\rm d}x\leq e^{C\;t}\int_{\mathbb{R}^{d}}\rho_{{\rm in}}(1+|x|^{2})\,{\rm d}x.$
  4. (iv)
    Entropy bound:
    $(\int_{\mathbb{R}^{d}}\rho|\log\rho|\,{\rm d}x)(T)+\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\lvert\nabla\rho_{\varepsilon}\rvert^{2}}{\rho_{\varepsilon}}\,{\rm d}x\,{\rm d}t\leq\bar{C}[\rho_{{\rm in}}].$
  5. (iv)
    $L^{\infty}$ smoothing:
    $\sup_{t>0}\|\rho(t)\|_{L^{\infty}(\mathbb{R}^{d})}\leq C\|\rho_{{\rm in}}\|_{L^{p}(\mathbb{R}^{d})}\left(1+\frac{1}{t^{\frac{d}{2p}}}\right).$
  6. (v)
    $L^{p}$ bound of the gradient:
    $\|\nabla\rho_{\varepsilon}\|_{L^{p}((t,T)\times\mathbb{R}^{d})}<C(\rho_{{\rm in}}),\quad\forall\;t>0.$
The proof contains two approximation steps: initially we introduce viscosity and regularize the initial data $f_{in}$ with a sequence of very smooth functions that converge (strongly in $L^{p}$) to $f_{in}$. We establish the existence of solutions to the approximated problem and subsequently remove the viscosity. In the last step we show how the solution obtains the initial data strongly in $L^{p}$ sense.
The following section provides a concise summary of the most relevant a-priori estimates. Section 3 delves into the global well-posedness of the model with viscosity, while Section 4 focuses on the vanishing viscosity limit and establishes uniqueness.
2. A priori estimates and technical results
In this section we give some a priori estimates of the solutions to
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\partial_{t}\rho\varphi dxdt=-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}\nabla\rho_{\varepsilon}\nabla\varphi\;dxdt$ (11)
[…]
for any $\varphi\in L^{2}(0,T,H^{1}(\mathbb{R}^{d}))$ with initial smooth data $\rho_{0}$. The function $\rho_{0}$ belongs to $W^{1,p}\cap L^{\infty}\cap L^{1}_{2}(\mathbb{R}^{d})$, with $p>d$.
The following a priori estimates hold:
Lemma 1 (Mass evolution).
Let $\rho_{\varepsilon}$ be as stated, then
$\int_{\mathbb{R}^{d}}\rho_{\varepsilon}(t)=e^{-\varepsilon t}\int_{\mathbb{R}^{d}}\rho_{0}.$
Proof.
Testing (11) with the test function $\varphi=\frac{1}{\left(1+\frac{|x|^{2}}{R^{2}}\right)^{\alpha/2}}$, with some $\alpha>{d}$, we get
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x)(T)=$ $\;(1+\varepsilon)\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\Delta\varphi\,{\rm d}x\,{\rm d}t+\chi\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\frac{\nabla\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\nabla\varphi\,{\rm d}x\,{\rm d}t$ (12)
$\,-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{-}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t+\int_{\mathbb{R}^{d}}\rho_{0}\varphi\,{\rm d}x.$
But notice that
-- $\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{-}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t=\int_{0}^{T}\int_{\mathbb{R}^{d}\cap\{|\nabla\rho_{\varepsilon}|\leq\chi\rho_{\varepsilon}\}}\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t$
$\leq\int_{0}^{T}\int_{\mathbb{R}^{d}\cap\{|\nabla\rho_{\varepsilon}|\leq\chi\rho_{\varepsilon}\}}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t-\chi\int_{0}^{T}\int_{\mathbb{R}^{d}\cap\{|\nabla\rho_{\varepsilon}|\leq\chi\rho_{\varepsilon}\}}\rho_{\varepsilon}\frac{\nabla\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\nabla\varphi\,{\rm d}x\,{\rm d}t$
$\leq{(1+\chi)}\int_{0}^{T}\int_{\mathbb{R}^{d}\cap\{|\nabla\rho_{\varepsilon}|\leq\chi\rho_{\varepsilon}\}}\rho_{\varepsilon}|\nabla\varphi|\,{\rm d}x\,{\rm d}t.$
So, introducing the estimate above in (12) we get
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x)(T)\leq$ $(1+\varepsilon)\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\Delta\varphi\,{\rm d}x\,{\rm d}t+(2\chi+1)\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left|\nabla\varphi\right|\,{\rm d}x\,{\rm d}t+\int_{\mathbb{R}^{d}}\rho_{0}\varphi\,{\rm d}x$
$=$ $\;(1+\varepsilon)\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left[\frac{\alpha(\frac{\alpha}{2}+1)\frac{|x|^{2}}{R^{4}}}{\left(1+\frac{|x|^{2}}{R^{2}}\right)^{\frac{\alpha}{2}+2}}-\frac{\alpha d\frac{1}{R^{2}}}{\left(1+\frac{|x|^{2}}{R^{2}}\right)^{\frac{\alpha}{2}+1}}\right]\,{\rm d}x\,{\rm d}t$
$+3\chi\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}\alpha\frac{|x|}{R^{2}}}{\left(1+\frac{|x|^{2}}{R^{2}}\right)^{\frac{\alpha}{2}+1}}\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t+\int_{\mathbb{R}^{d}}\rho_{0}\varphi\,{\rm d}x$
$\leq$ $\;(1+\varepsilon)\frac{\alpha}{2R}\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}(\frac{\alpha}{2}+1-d)}{\left(1+\frac{|x|^{2}}{R^{2}}\right)^{\frac{\alpha}{2}+1}}\,{\rm d}x\,{\rm d}t+3\chi\frac{\alpha}{R}\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}}{\left(1+\frac{|x|^{2}}{R^{2}}\right)^{\frac{\alpha}{2}}}$
$-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t+\int_{\mathbb{R}^{d}}\rho_{0}\varphi\,{\rm d}x.$
Now, we choose $\alpha$ such that $\frac{\alpha}{2}+1-d>0$ and get
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x)(T)\leq\;\frac{C}{R}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t+\int_{\mathbb{R}^{d}}\rho_{0}\varphi\,{\rm d}x,$ (13)
where $C>0$. Gronwall’s Lemma yields
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x)(T)\leq e^{\frac{CT}{R}}\int_{\mathbb{R}^{d}}\rho_{0}\varphi\,{\rm d}x\leq e^{\frac{CT}{R}}\int_{\mathbb{R}^{d}}\rho_{0}\,{\rm d}x.$
By monotone convergence we can take the limit $R\rightarrow\infty$ and conclude that $\int\rho_{\varepsilon}\leq\int\rho_{0}$. Now that we know that the mass of $\rho_{\varepsilon}$ is bounded, we test again (11) with $\varphi$ , but this time we pass directly to the limit $R\rightarrow\infty$ in (12) to find
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x)(T)=-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x+\int_{\mathbb{R}^{d}}\rho_{0}\,{\rm d}x.$ (14)
The above identity can be rewritten as an initial value problem for an ODE, which can be easily solved, yielding
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x)(T)=e^{-\varepsilon T}\int_{\mathbb{R}^{d}}\rho_{0}\,{\rm d}x,$
which finishes the proof. ∎
Proposition 1 ($L^{p}$ estimates).
Let $\rho_{\varepsilon}$ be as stated, then
$\|\rho_{\varepsilon}\|_{L^{\infty}((0,T);L^{p}(\mathbb{R}^{d}))}\leq\|\rho_{0}\|_{L^{p}(\mathbb{R}^{d})},\qquad\|\nabla\rho_{\varepsilon}^{\frac{p}{2}}\|_{L^{2}((0,T)\times\mathbb{R}^{d})}\leq C,$
where $C$ only depends on $\|\rho_{0}\|_{L^{p}},p$ and $\chi$.
Proof.
First, notice that testing (11) with $\rho_{\varepsilon}^{p-1}$we get that $\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{p}(t)\leq\int_{\mathbb{R}^{d}}\rho_{0}^{p}$. Then, we compute
$\frac{1}{p}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{p}(t)\leq$ $-(p-1)\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{p-2}\lvert\nabla\rho_{\varepsilon}\rvert^{2}+\chi(p-1)\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{p-1}\lvert\nabla\rho_{\varepsilon}\rvert+\frac{1}{p}\int_{\mathbb{R}^{d}}\rho_{0}^{p}$
$\leq$ $-\frac{(p-1)}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{p-2}\lvert\nabla\rho_{\varepsilon}\rvert^{2}+\chi^{2}(p-1)\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{p}+\frac{1}{p}\int_{\mathbb{R}^{d}}\rho_{0}^{p}$
$\leq$ $-\frac{2(p-1)}{p^{2}}\int_{0}^{T}\int_{\mathbb{R}^{d}}\lvert\nabla\rho_{\varepsilon}^{\frac{p}{2}}\rvert^{2}+\chi^{2}(p-1)\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{0}^{p}+\frac{1}{p}\int_{\mathbb{R}^{d}}\rho_{0}^{p}.$
We get that
$\|\rho_{\varepsilon}\|^{p}_{L^{\infty}(0,T;L^{p}(\mathbb{R}^{d}))}+\|\nabla\rho_{\varepsilon}^{\frac{p}{2}}\|_{L^{2}(0,T;L^{2}(\mathbb{R}^{d}))}\leq C,$
with the right hand side independent of $\varepsilon$. ∎
Proposition 2 (Moment Evolution).
Let $\rho_{\varepsilon}$ be as stated, then
$\int_{\mathbb{R}^{d}}\rho_{\varepsilon}(t)(1+|x|^{2})\,{\rm d}x\leq e^{C\;t}\int_{\mathbb{R}^{d}}\rho_{0}(1+|x|^{2})\,{\rm d}x,$
where $C$ is independent of $\varepsilon$.
Proof.
We test the equation with $(1+|x|^{2})\varphi$, where $\varphi=\frac{1}{\left(1+\frac{|x|^{2}}{R^{2}}\right)^{\alpha/2}}$ as in Lemma 1. Then, by Young’s inequality, noticing that $|\nabla\varphi|,|\Delta\varphi|\leq\varphi$ and $0\leq\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{-}|\nabla\rho_{\varepsilon}|\leq 2\chi\rho_{\varepsilon}$, we get
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}(1+|x|^{2})\varphi\,{\rm d}x)(T)\leq C\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}(1+|x|^{2})\varphi\,{\rm d}x\,{\rm d}t+\int_{\mathbb{R}^{d}}\rho_{0}(1+|x|^{2})\varphi\,{\rm d}x.$
Gronwall’s lemma gives
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}(1+|x|^{2})\varphi\,{\rm d}x)(T)\leq e^{C\;t}\int_{\mathbb{R}^{d}}\rho_{0}(1+|x|^{2})\varphi\,{\rm d}x.$
By dominated convergence we pass to the limit in $R$ and get the result. ∎
Proposition 3 ( Estimate for $\|\nabla\sqrt{\rho_{\varepsilon}}\|_{L^{2}((0,T)\times\mathbb{R}^{d})}$).
Let $\rho_{\varepsilon}$ be as stated, then
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\log\rho_{\varepsilon}\,{\rm d}x)(T)+\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\lvert\nabla\rho_{\varepsilon}\rvert^{2}}{\rho_{\varepsilon}}\,{\rm d}x\,{\rm d}t\leq\bar{C}[\rho_{0}]$
where $\bar{C}$ depends on the moment, entropy and $L^{2}$ norm of the initial data.
Proof.
Let us test (11) with the test function $\varphi=\log(\rho_{\varepsilon}+\sigma)-\log{\sigma}$, with $10<\sigma<<$ 1. After some computations we get
$\left(\int_{\mathbb{R}^{d}}\left[(\rho_{\varepsilon}+\sigma)\log(\rho_{\varepsilon}+\sigma)-\rho_{\varepsilon}\log\sigma\right]\,{\rm d}x\right)(T)+\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\left[\log(\rho_{\varepsilon}+\sigma)-\log{\sigma}\right]\,{\rm d}x\,{\rm d}t$
$\qquad\leq\int_{\mathbb{R}^{d}}\left[(\rho_{0}+\sigma)\log(\rho_{0}+\sigma)-\rho_{0}\log\sigma\right]\,{\rm d}x,$
which implies
$\left(\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\log(\rho_{\varepsilon}+\sigma)\,{\rm d}x\right)(T)\leq$ $\int_{\mathbb{R}^{d}}(\rho_{0}+\sigma)\log(\rho_{0}+\sigma)\,{\rm d}x$
$-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\log{(\rho_{\varepsilon}+\sigma)}\,{\rm d}x\,{\rm d}t$
$+\log\sigma\left(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x-\int_{\mathbb{R}^{d}}\rho_{0}\,{\rm d}x+\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x\,{\rm d}t\right).$
The last term in the previous inequality is zero by (14). Taking $\sigma\rightarrow 0$, by monotone and dominated convergence, we get
$\left(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\log\rho_{\varepsilon}\,{\rm d}x\right)(T)\leq$ $\int_{\mathbb{R}^{d}}\rho_{0}\log\rho_{0}\,{\rm d}x-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\log{\rho_{\varepsilon}}\,{\rm d}x\,{\rm d}t.$
In particular,
$\left(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}|\log\rho_{\varepsilon}|\,{\rm d}x\right)(T)\leq e^{\varepsilon T}C[\rho_{0}],$
where $C[\rho_{0}]$ is a constant that depends on $\int_{\mathbb{R}^{d}}\rho_{0}|\log{\rho_{0}}|\,{\rm d}x$, the mass and the second moment of $\rho_{0}$. To bound $\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\varepsilon}|^{2}}{\rho_{\varepsilon}}$, we define $h[\rho_{\varepsilon}](t):=\left(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\log(\rho_{\varepsilon})\,{\rm d}x\right)(t)$. Using the fact that
$\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}=\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)+\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{-},$
and repeating the same computations as before, we obtain
$h[\rho_{\varepsilon}](T)+\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\varepsilon}|^{2}}{\rho_{\varepsilon}+\sigma}\,{\rm d}x\,{\rm d}t\leq\frac{\chi^{2}}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}h[\rho_{\varepsilon}](t)\,{\rm d}t+h[\rho_{0}],$
which implies
$\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\varepsilon}|^{2}}{\rho_{\varepsilon}+\sigma}\,{\rm d}x\,{\rm d}t\leq\frac{\chi^{2}}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\,{\rm d}x\,{\rm d}t+e^{\varepsilon T}C[\rho_{0}].$
Lemma 2 ($\|\nabla\rho_{\varepsilon}\|_{L^{p}}$ estimate).
Let $\rho_{\varepsilon}$ be as stated, then
$\|\nabla\rho_{\varepsilon}\|_{L^{p}((t,T)\times\mathbb{R}^{d})}<\frac{1}{t^{1/2}}C(\|\rho_{0}\|_{L^{p}}),\quad\forall\;t>0.$
Proof.
Define $f$ as
$f:=\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}-\nabla\rho_{\varepsilon}.$
Note that each component of $f$ belongs to $L^{\infty}(0,T,L^{q}(\mathbb{R}^{d}))$ for any $q\in[1,+\infty)$. In fact, in the domain where $|\nabla\rho_{\varepsilon}|\leq\chi\rho_{\varepsilon}$, the vector function $f$ coincides with $\nabla\rho_{\varepsilon}$, which is bounded by $\chi\rho_{\varepsilon}$. In the complementary domain, we have that
$\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}=\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right),$
and $|f|\leq\chi\rho_{\varepsilon}$. With this in mind, consider the function $u_{i}$, solution to
${u_{i}}_{t}-(1+\varepsilon)\Delta u_{i}+\varepsilon u_{i}=f_{i},\quad u_{i}(x,0)=0,\quad i=1,...,d.$
Parabolic regularity theory tells us that $u_{i}\in L^{q}(0,T,W^{2,q})$, with bounds uniformly in $\varepsilon$. Direct computations shows that $\rho_{\varepsilon}$ can be written as the sum of $\textrm{div}(u)+g$, where $g$ solves
$g_{t}=(1+\varepsilon)\Delta g-\varepsilon g,\quad g(x,0)=\rho_{0}.$
Since $\rho_{0}\in L^{p}(\mathbb{R}^{d})$, we have that
$\sup_{t>0}\|\nabla g\|_{L^{p}}\leq\frac{c}{t^{1/2}}\|\rho_{0}\|_{L^{p}},$
with $c$ independent on $\varepsilon$. Summarizing, $\textrm{div}(u)\in L^{p}(0,T,W^{1,p})$, $g\in L^{\infty}(t,T,W^{1,p})$, and the thesis follows. ∎
Lemma 3.
Let $\rho_{\varepsilon}$ be as stated, then for every $\tau>0$ it holds
$\|\rho_{\varepsilon}\|_{L^{\infty}({\tau},T;\mathbb{R}^{d})}\leq C{T^{\frac{1}{p}}\left(1+\tau^{-\frac{d+2}{2p}}\right)}\|\rho_{0}\|_{L^{p}(\mathbb{R}^{d})},$
where $C$ only depends on $p$, $\chi$ and other universal constants.
Proof.
Let us consider the following sequence of times
$T_{k}=\tau(1-2^{-k}),$
and the following sequence of energies
$U_{k}=\int_{T_{k}}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k}^{p}\,{\rm d}x\,{\rm d}t,\quad\mbox{where}\quad\rho_{\varepsilon,k}=(\rho_{\varepsilon}-C_{k})_{+},\quad\mbox{and}\quad C_{k}=M(1-2^{-k}),$
with $M>0$ to fix later. Then, we test (11) with $\rho_{\varepsilon,k+1}^{p-1}$ and we integrate in time in $(t_{1},t_{2})$, where $0\leq T_{k}\leq t_{1}\leq T_{k+1}\leq t_{2}\leq T$. We find
$\frac{1}{p}\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{d}}$ $\partial_{t}(\rho_{\varepsilon,k+1}^{p})\,{\rm d}x\,{\rm d}t$
$\leq-\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{d}}\nabla\rho_{\varepsilon}\nabla(\rho_{\varepsilon,k+1}^{p-1})\,{\rm d}x\,{\rm d}t+\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{d}}\frac{\chi\rho}{|\nabla\rho_{\varepsilon}|}\nabla(\rho_{\varepsilon,k+1}^{p-1})\nabla\rho_{\varepsilon}\,{\rm d}x\,{\rm d}t.$
After some computations, from the expression above we get
$\frac{1}{p}\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{d}}\partial_{t}(\rho_{\varepsilon,k+1}^{p})\,{\rm d}x\,{\rm d}t+\frac{(p-1)}{p^{2}}\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{d}}|\nabla\rho_{\varepsilon,k+1}^{\frac{p}{2}}|^{2}\,{\rm d}x\,{\rm d}t\leq\frac{(p-1)}{2}\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k+1}^{p}\,{\rm d}x\,{\rm d}t$
$+2\chi^{2}C_{k+1}^{2}(p-1)\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k+1}^{p-2}\,{\rm d}x\,{\rm d}t.$ (15)
Now, we notice that, in the set $\rho_{\varepsilon,k+1}>0$, $\rho_{\varepsilon,k}=\rho_{\varepsilon,k+1}+C_{k+1}-C_{k}$, i.e. that $\frac{\rho_{\varepsilon,k}}{C_{k+1}-C_{k}}=\frac{\rho_{\varepsilon,k+1}}{C_{k+1}-C_{k}}+1>1$. Then, for any $\alpha>0$ we have $\left(\frac{\rho_{\varepsilon,k}}{C_{k+1}-C_{k}}\right)^{\alpha}>1$ and, consequently $\left(\frac{1}{C_{k+1}-C_{k}}\right)^{\alpha}\rho_{\varepsilon,k}^{1+\alpha}>\rho_{\varepsilon,k+1}$. Now, let us choose $\alpha=\frac{2}{p-2}$ to get
$\rho_{\varepsilon,k+1}\leq\left(\frac{1}{C_{k+1}-C_{k}}\right)^{\frac{2}{p-2}}\rho_{\varepsilon,k}^{\frac{p}{p-2}}\mathds{1}_{\{\rho_{\varepsilon,k+1}>0\}}.$
Introducing this expression in (2) we get
$\frac{1}{p}\left(\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k+1}^{p}\,{\rm d}x\right)(t_{2})+\frac{(p-1)}{p^{2}}\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{d}}|\nabla\rho_{\varepsilon,k+1}^{\frac{p}{2}}|^{2}\,{\rm d}x\,{\rm d}t\leq\frac{(p-1)}{2}\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k+1}^{p}\,{\rm d}x\,{\rm d}t$
$+2\chi^{2}(p-1)\left(\frac{C_{k+1}}{C_{k+1}-C_{k}}\right)^{2}\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k}^{p}\,{\rm d}x\,{\rm d}t+\frac{1}{p}\left(\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k+1}^{p}\,{\rm d}x\right)(t_{1}).$
Then, averaging $t_{1}$ in $(T_{k},T_{k+1})$ and taking the supremum in $t_{2}\in(T_{k+1},T)$ we have
$E_{k+1}$ $:=\sup_{t\in[T_{k+1},T]}\left(\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k+1}^{p}\,{\rm d}x\right)+\int_{T_{k+1}}^{T}\int_{\mathbb{R}^{d}}|\nabla\rho_{\varepsilon,k+1}^{\frac{p}{2}}|^{2}\,{\rm d}x\,{\rm d}t$
$\leq C\left(1+2^{k}+\frac{2^{k}}{{\tau}}\right)\int_{T_{k}}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k}^{p}\,{\rm d}x\,{\rm d}t,$
where $C$ only depends on $p$ and $\chi$. From now on, this $C$ will change from line to line for convenience, always depending only on $p$ and $\chi$. Now, notice that $E_{k+1}$ controls $\rho_{\varepsilon,k+1}$ in $L^{\infty}(T_{k+1},T;L^{p}(\mathbb{R}^{d}))$ and, thanks to Sobolev, $L^{p}(T_{k+1},T;L^{\frac{dp}{d-2}}(\mathbb{R}^{d}))$, so, by interpolation
$\|\rho_{\varepsilon,k+1}\|^{p}_{L^{\frac{p(d+2)}{d}}(\mathbb{R}^{d}\times(T_{k+1},T))}\leq CE_{k+1}\leq C^{\prime}\left(1+2^{k}+\frac{2^{k}}{{\tau}}\right)\int_{T_{k}}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k}^{p}\,{\rm d}x\,{\rm d}t.$ (16)
Then, Chebyshev’s inequality yields:
$U_{k+1}$ $:=\int_{T_{k+1}}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k+1}^{p}\,{\rm d}x\,{\rm d}t$
$\leq\left(\int_{T_{k+1}}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k+1}^{\frac{p(d+2)}{d}}\,{\rm d}x\,{\rm d}t\right)^{\frac{d}{d+2}}\left(\int_{T_{k+1}}^{T}\int_{\mathbb{R}^{d}}\mathds{1}_{\{\rho_{\varepsilon,k+1}>0\}}\,{\rm d}x\,{\rm d}t\right)^{\frac{2}{d+2}}$
$=\left(\int_{T_{k+1}}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k+1}^{\frac{p(d+2)}{d}}\,{\rm d}x\,{\rm d}t\right)^{\frac{d}{d+2}}\left(\int_{T_{k+1}}^{T}\int_{\mathbb{R}^{d}}\mathds{1}_{\{\rho_{\varepsilon,k}>\frac{M}{2^{k+1}}\}}\,{\rm d}x\,{\rm d}t\right)^{\frac{2}{d+2}}$
$\leq CE_{k+1}\left(\frac{2^{k+1}}{M}\right)^{\frac{2p}{d+2}}\left(\int_{T_{k}}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon,k}^{p}\,{\rm d}x\,{\rm d}t\right)^{\frac{2}{d+2}}$
$\leq\left(1{+\frac{1}{\tau}}\right)\frac{C^{k}}{M^{\frac{2p}{d+2}}}U_{k}^{\beta},$
where $\beta=1+\frac{2}{d+2}$, using (16) to obtain the second and the third inequality. From [9, Lemma 7.1] it follows that $U_{k}\rightarrow 0$ as $k\rightarrow\infty$, i.e. that $\rho_{\varepsilon}\leq M$, if
$\left(1+\frac{1}{{\tau}}\right)^{\frac{d+2}{2}}\frac{C}{M^{p}}U_{0}<1.$
So, we fix $M$ in order to
$C\left(1+\frac{1}{{\tau}}\right)^{\frac{d+2}{2p}}U_{0}^{1/p}\sim M.$
Finally, we conclude from Proposition 1
$\|\rho_{\varepsilon}\|_{L^{\infty}(0,T;\mathbb{R}^{d})}\leq M\sim C\left(1+\frac{1}{{\tau}}\right)^{\frac{d+2}{2p}}\|\rho_{\varepsilon}\|_{L^{p}(0,T;\mathbb{R}^{d})}\leq C{T^{\frac{1}{p}}}\left(1+\frac{1}{{\tau}^{\frac{d+2}{2p}}}\right)\|\rho_{0}\|_{L^{p}(\mathbb{R}^{d})}.$
Lemma 4 (Compactness Lemma).
Let $\{f_{n}\}$ be a bounded sequence in $L^{2}(0,T;H^{1}(\mathbb{R}^{d}))\cap L^{2}(0,T;\mathbb{R}^{d},\sqrt{1+|x|^{2}}\,{\rm d}x)$ such that $\partial_{t}f_{n}$ is bounded in $L^{2}(0,T;H^{-1}(\mathbb{R}^{d}))$, then
$f_{n}\rightarrow f\qquad\mbox{strongly in }\;L^{2}(0,T;\mathbb{R}^{d}),\qquad\mbox{as }n\rightarrow\infty.$
Proof.
As $\{f_{n}\}$ is uniformly bounded in $L^{2}(0,T;H^{1}(\mathbb{R}^{d}))$ we have that
$f_{n}\rightharpoonup f\quad\mbox{weakly in }L^{2}(0,T;L^{\frac{2d}{d-2}}(\mathbb{R}^{d})).$
Then, for any $R>0$, as $\partial_{t}f_{n}\in L^{2}(0,T;H^{-1}({B_{R}}))$ and $H^{1}(B_{R})\subset\subset L^{q}(B_{R})$ with $2\leq q<{\frac{2d}{d-2}}$, by Aubin-Lions Theorem (Theorem 5 in [16]), we have that a subsequence, $\{f_{n}^{R}\}$, converges strongly to $f$ as follows
$f_{n}^{R}\rightarrow f\quad\mbox{strongly in }L^{2}(0,T;L^{q}(B_{R})).$
By a Cantor diagonal argument, we have a new subsequence $f_{n}$, independent of $R$, such that
$f_{n}\rightarrow f\quad\mbox{strongly in }L^{2}(0,T;L^{q}(B_{R})),$
and consequently,
$f_{n}\rightarrow f\quad\mbox{a.e. in }[0,T]\times\mathbb{R}^{d}.$
By Fatou’s Lemma, we also have
$\int_{0}^{T}\int_{\mathbb{R}^{d}}f^{2}\sqrt{1+|x|^{2}}\,{\rm d}x\,{\rm d}t\leq\liminf\limits_{\varepsilon\rightarrow 0}\int_{0}^{T}\int_{\mathbb{R}^{d}}f_{n}^{2}\sqrt{1+|x|^{2}}\,{\rm d}x\,{\rm d}t<\infty.$
Now, setting $q=2$, for any $\delta>0$, there exist sufficiently large $n$ and $R$ such that
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\lvert f_{n}-f\rvert^{2}\,{\rm d}x\,{\rm d}t=\int_{0}^{T}\int_{B_{R}}\lvert f_{n}-f\rvert^{2}\,{\rm d}x\,{\rm d}t+\int_{0}^{T}\int_{\mathbb{R}^{d}\setminus B_{R}}\lvert f_{n}-f\rvert^{2}\,{\rm d}x\,{\rm d}t$
$\leq\frac{\delta}{2}+\frac{1}{\sqrt{1+R^{2}}}\int_{0}^{T}\int_{\mathbb{R}^{d}\setminus B_{R}}\lvert f_{n}-f\rvert^{2}\sqrt{1+|x|^{2}}\,{\rm d}x\,{\rm d}t\leq\delta,$
and the thesis follows. ∎
3. Existence for the approximated equation
This section is devoted to show existence of solutions to (11) with smooth initial data. First, we semi-linearize (11) and obtain a monotone functional inside the divergence operator. Then, by standard theory for monotone operators we prove existence of this new problem. A Leray-Schauder fixed point will yield existence of solutions to (11).
3.1. Existence for the semi-linearized equation
For given $\varepsilon$, $\sigma\in[0,1]$ and $z\in L^{2}(0,T;L^{2}(\mathbb{R}^{d}))$, let us present the following regularized weak formulation of (9)
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\partial_{t}\rho_{\varepsilon}\,\varphi\,{\rm d}x\,{\rm d}t=$ $-\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi z_{+}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t$
$-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t,$ (17)
$\forall\varphi\in L^{2}(0,T;H^{1}(\mathbb{R}^{d}))$, where $\rho_{\varepsilon}(0,\cdot)=\rho_{{\rm in},\varepsilon}$ and $\rho_{{\rm in},\varepsilon}$ is a $W^{1,p}\cap L^{\infty}(\mathbb{R}^{d})$ approximation of $\rho_{{\rm in}}$ such that $\rho_{{\rm in},\varepsilon}\rightarrow\rho_{{\rm in}}$ strongly in $L^{p}(\mathbb{R}^{d})$ when $\varepsilon\rightarrow 0$. In order to prove the existence of the above equation, we use general theory of monotone operators, concretely, [18, Theorem 30.A]. We define the following operator $A$ as
$\langle A[u](t),\varphi\rangle=$ $\sigma\int_{\mathbb{R}^{d}}\left(1-\frac{\chi z_{+}}{\lvert\nabla u(t,x)\rvert}\right)_{+}\nabla u(t,x)\nabla\varphi\,{\rm d}x$
$+\varepsilon\int_{\mathbb{R}^{d}}\nabla u(t,x)\nabla\varphi\,{\rm d}x+\varepsilon\int_{\mathbb{R}^{d}}u(t,x)\varphi\,{\rm d}x,$
which must satisfy the following properties: 1. Monotonicity. . The operator $A$ is monotone, since $\forall\;u,v\in H^{1}(\mathbb{R}^{d})$ we have
$\langle A[u]-A[v],u-v\rangle=$ $\sigma\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi z_{+}}{\lvert\nabla u\rvert}\right)_{+}\nabla u-\left(1-\frac{\chi z_{+}}{\lvert\nabla v\rvert}\right)_{+}\nabla v\right]\nabla(u-v)\,{\rm d}x$
$+\varepsilon\|u-v\|^{2}_{H^{1}(\mathbb{R}^{d})}\geq 0.$
2. Hemicontinuity. Let us check that the map
$t\mapsto\langle A[u+tv],\varphi\rangle$
is continuous in $t\in[0,1]$ for all $u,v,\varphi\in H^{1}(\mathbb{R}^{d})$. Notice that
$\left\lvert\left(1-\frac{\chi z_{+}}{\lvert\nabla(u+tv)\rvert}\right)_{+}\nabla(u+tv)\nabla\varphi\right\rvert\leq\lvert\nabla u\rvert\lvert\nabla\varphi\rvert+\lvert\nabla v\rvert\lvert\nabla\varphi\rvert$
so
$\lim_{t\rightarrow t_{0}}\langle A[u+tv],\varphi\rangle=$ $\sigma\int_{\mathbb{R}^{d}}\left(1-\frac{\chi z_{+}}{\lvert\nabla(u+t_{0}v)\rvert}\right)_{+}\nabla(u+t_{0}v)\nabla\varphi\,{\rm d}x+\varepsilon\int_{\mathbb{R}^{d}}\nabla(u+t_{0}v)\nabla\varphi\,{\rm d}x$
$+\varepsilon\int_{\mathbb{R}^{d}}(u+t_{0}v)\varphi\,{\rm d}x=\langle A(u+t_{0}v),\varphi\rangle,$
by dominated convergence. 3. Coerciveness. We just compute
$\langle A[u],u\rangle=\sigma\int_{\mathbb{R}^{d}}\left(1-\frac{\chi z_{+}}{\lvert\nabla u\rvert}\right)_{+}\lvert\nabla u\rvert^{2}\,{\rm d}x+\varepsilon\|u\|^{2}_{H^{1}(\mathbb{R}^{d})}\geq\varepsilon\|u\|_{H(\mathbb{R}^{d})}^{2},$
so $A$ is coercive. 4. Growth condition. For this we have to prove that $\|A[u]\|_{H^{*}(\mathbb{R}^{d})}\leq C\|u\|_{H(\mathbb{R}^{d})}$, but, by Hölder,
$\|A[u]\|_{H^{*}(\mathbb{R}^{d})}=$ $\sup_{\begin{subarray}{c}\varphi\in H^{1}(\mathbb{R}^{d})\\ \|\varphi\|\leq 1\end{subarray}}\langle A[u],\varphi\rangle$
$=$ $\sup_{\begin{subarray}{c}\varphi\in H^{1}(\mathbb{R}^{d})\\ \|\varphi\|\leq 1\end{subarray}}\int_{\mathbb{R}^{d}}\left[\sigma\left(1-\frac{\chi z_{+}}{\lvert\nabla u\rvert}\right)_{+}\nabla u\nabla\varphi+\varepsilon\nabla u\nabla\varphi+\varepsilon u\varphi\right]\,{\rm d}x$
$\leq$ $C\|u\|_{H(\mathbb{R}^{d})}.$
5. Measurability. By assumption. Then, by [18, Theorem 30.A], there exist a unique weak solution $\rho_{\varepsilon}\in L^{2}(0,T;H^{1}(\mathbb{R}^{d}))$ to (3.1) with $\partial_{t}\rho\in L^{2}(0,T;H^{-1}(\mathbb{R}^{d}))$ satisfying the initial condition in the sense of the strong $L^{2}(\mathbb{R}^{d})$ convergence:
$\rho_{\varepsilon}(t)\to\rho_{in,\varepsilon}\quad\mbox{strongly in $L^{2}(\mathbb{R}^{d})$ as $t\to 0$.}$ (18)
Therefore we can define the mapping
$\mathcal{T}:(z,\sigma)\in L^{2}((0,T)\times\mathbb{R}^{d})\times[0,1]\mapsto\rho_{\varepsilon}\in L^{2}((0,T)\times\mathbb{R}^{d}),$
where $\rho_{\varepsilon}$ is the unique solution to (3.1).
3.2. Existence for approximated solutions
Now, we are going to use the Leray-Schauder fixed point theorem to show the existence of a solution to (11). To use the theorem we require $\mathcal{T}$ to be compact. First, using $\rho_{\varepsilon}$ as test function in (3.1) we get
$\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\partial_{t}(\rho_{\varepsilon}^{2})\,{\rm d}x\,{\rm d}t=$ $-\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi(z_{+}+\varepsilon)}{\lvert\nabla\rho(t,x)\rvert}\right)_{+}|\nabla\rho_{\varepsilon}|^{2}\,{\rm d}x\,{\rm d}t$
$-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}|\nabla\rho_{\varepsilon}|^{2}\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{2}\,{\rm d}x\,{\rm d}t,$
which implies:
[…]
Thus, the $L^{2}(0,T;H^{1}(\mathbb{R}^{d}))$ norm of any solution $\rho_{\varepsilon}$ has a uniform bound. Now, we test (3.1) with $\varphi=\rho_{\varepsilon}\sqrt{1+|x|^{2}}\phi_{R}$, where $\phi_{R}\in C^{\infty}_{c}(B_{2R})$ is such that $\phi_{R}\equiv 1$ in $B_{R}$ and $|\nabla\phi_{R}|\leq\frac{C}{R}$ for some $C>0$. Doing this we get
$\frac{1}{2}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{2}\sqrt{1+\lvert x\rvert^{2}}\phi_{R}\,{\rm d}x\leq$ $\;\;\frac{1}{2}\int_{\mathbb{R}^{d}}\rho_{{\rm in},\varepsilon}^{2}\sqrt{1+\lvert x\rvert^{2}}\,{\rm d}x+(\sigma+\varepsilon)\int_{0}^{T}\int_{B_{2R}}\frac{\rho_{\varepsilon}|\nabla\rho_{\varepsilon}|\cdot|x|}{\sqrt{1+\lvert x\rvert^{2}}}\,{\rm d}x\,{\rm d}t$
$+(\sigma+\varepsilon)\int_{0}^{T}\int_{B_{2R}}|\nabla\rho_{\varepsilon}|\rho_{\varepsilon}\sqrt{1+|x|^{2}}|\nabla\phi_{R}|\,{\rm d}x\,{\rm d}t.$
Then, using $\frac{\lvert x\rvert}{\sqrt{1+\lvert x\rvert^{2}}}\leq 1$, Young’s inequality and that $|\nabla\phi_{R}|\leq\frac{C}{R}$, we get
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}^{2}\sqrt{1+\lvert x\rvert^{2}}\phi_{R}\,{\rm d}x\,{\rm d}t\leq$ $\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{{\rm in},\varepsilon}^{2}\sqrt{1+\lvert x\rvert^{2}}\,{\rm d}x\,{\rm d}t+c\|\rho_{\varepsilon}\|^{2}_{L^{2}(0,T;H^{1}(\mathbb{R}^{d}))}$
$+\frac{C\sqrt{1+4|R|^{2}}}{R}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}|\nabla\rho_{\varepsilon}|\,{\rm d}x\,{\rm d}t$
$\leq$ $C(\rho_{{\rm in},\varepsilon}).$
Taking the limit $R\rightarrow\infty$ we get that solutions to (3.1) are bounded in $L^{2}(0,T;H^{1}(\mathbb{R}^{d}))\cap L^{2}(0,T;\mathbb{R}^{d},\sqrt{1+|x|^{2}}\,{\rm d}x)$. Also, for any $\varphi\in L^{2}(0,T,H^{1}(\mathbb{R}^{d}))$ we have
$\left\lvert\int_{0}^{T}\int_{\mathbb{R}^{d}}\partial_{t}\rho_{\varepsilon}(t)\varphi\,{\rm d}x\,{\rm d}t\right\rvert\leq C\|\rho_{\varepsilon}(t)\|_{H^{1}(\mathbb{R}^{d})}\|\varphi\|_{L^{2}(0,T,H^{1}(\mathbb{R}^{d}))}\leq C\|\varphi\|_{L^{2}(0,T,H^{1}(\mathbb{R}^{d}))}.$ (19)
By Lemma 4, $\mathcal{T}$ maps bounded sets to precompact sets. Now, we have to prove that the map is continuous, i.e. that $\lim\limits_{n\rightarrow\infty}\mathcal{T}(\sigma_{n},z_{n})=\mathcal{T}(\sigma,z)$ when $z_{n}\rightarrow z$ strongly in $L^{2}(0,T,L^{2}(\mathbb{R}^{d}))$ and $\sigma_{n}\rightarrow\sigma$ in $[0,1]$. Is easy to see that $\lim\limits_{n\rightarrow\infty}\mathcal{T}(\sigma_{n},z_{n})=\lim\limits_{n\rightarrow\infty}\mathcal{T}(\sigma,z_{n})$. Then, as the map is compact, it only remains to prove
$\lim_{n\rightarrow\infty}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,n}\rvert}\right)_{+}\nabla\rho_{\varepsilon,n}\nabla\varphi\,{\rm d}x\,{\rm d}t=\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi z_{+}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t,$
but as $\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,n}\rvert}\right)_{+}\nabla\rho_{\varepsilon,n}$ is uniformly bounded in $L^{2}(0,T,L^{2}(\mathbb{R}^{d}))$, it has a weakly convergent subsequence and we only need $\nabla\rho_{\varepsilon,n}\rightarrow\nabla\rho_{\varepsilon}$ a.e. $x\in\mathbb{R}^{d}$. To show this convergence, we prove that $\{\nabla\rho_{\varepsilon,k}\}$ is a Cauchy sequence in $L^{2}(0,T;L^{2})$: let $\rho_{\varepsilon,n}$ and $\rho_{\varepsilon,k}$ solutions to (3.1) with $z_{n}$ and $z_{k}$ respectively. We subtract the equation for $\rho_{\varepsilon,k}$ from the one of $\rho_{\varepsilon,n}$ and get, after some manipulations:
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\partial_{t}$ $(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\varphi\,{\rm d}x\,{\rm d}t$
$=$ $-\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,n}\rvert}\right)_{+}\nabla\rho_{\varepsilon,n}-\left(1-\frac{\chi(z_{k})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\nabla\rho_{\varepsilon,k}\right]\nabla\varphi\,{\rm d}x\,{\rm d}t$
$-\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\nabla\rho_{\varepsilon,k}-\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\nabla\rho_{\varepsilon,k}\right]\nabla\varphi\,{\rm d}x\,{\rm d}t$
$-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(\nabla\rho_{\varepsilon,n}-\nabla\rho_{\varepsilon,k}\right)\nabla\varphi\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(\rho_{\varepsilon,n}-\rho_{\varepsilon,k}\right)\varphi\,{\rm d}x\,{\rm d}t$
$=$ $-\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,n}\rvert}\right)_{+}\nabla\rho_{\varepsilon,n}-\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\nabla\rho_{\varepsilon,k}\right]\nabla\varphi\,{\rm d}x\,{\rm d}t$
$-\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}-\left(1-\frac{\chi(z_{k})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\right]\nabla\rho_{\varepsilon,k}\nabla\varphi\,{\rm d}x\,{\rm d}t$
$-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(\nabla\rho_{\varepsilon,n}-\nabla\rho_{\varepsilon,k}\right)_{+}\nabla\varphi\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(\rho_{\varepsilon,n}-\rho_{\varepsilon,k}\right)\varphi\,{\rm d}x\,{\rm d}t.$
We use the test function $\varphi=\rho_{\varepsilon,n}-\rho_{\varepsilon,k}$ and get
$\left(\int_{\mathbb{R}^{d}}\rvert\rho_{\varepsilon,n}-\rho_{\varepsilon,k}\lvert^{2}\,{\rm d}x\right)(t)$
$\leq$ $-\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,n}\rvert}\right)_{+}\nabla\rho_{\varepsilon,n}-\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\nabla\rho_{\varepsilon,k}\right]\nabla(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\,{\rm d}x\,{\rm d}t$
$-\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}-\left(1-\frac{\chi(z_{k})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\right]\nabla\rho_{\varepsilon,k}\cdot\nabla(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\,{\rm d}x\,{\rm d}t$
$-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\lvert\nabla(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\rvert^{2}\,{\rm d}x\,{\rm d}t$
$\leq$ $-\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,n}\rvert}\right)_{+}\nabla\rho_{\varepsilon,n}-\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\nabla\rho_{\varepsilon,k}\right]\nabla(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\,{\rm d}x\,{\rm d}t$
$+\frac{\sigma^{2}}{2\varepsilon}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}-\left(1-\frac{\chi(z_{k})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\right]^{2}|\nabla\rho_{\varepsilon,k}|^{2}$
$-\frac{\varepsilon}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\lvert\nabla(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\rvert^{2}\,{\rm d}x\,{\rm d}t,$ (20)
using the Young’s inequality in the second estimate. Now come two key points in the argument. The first one is that, due to monotonicity, we have
$-\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,n}\rvert}\right)_{+}\nabla\rho_{\varepsilon,n}-\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\nabla\rho_{\varepsilon,k}\right]\nabla(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\,{\rm d}x\,{\rm d}t\leq 0.$
The second one is that $\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}$ is Lipschitz continuous with respect to $(z_{n})_{+}$. In fact, we have the following
$\left|\left(1-\frac{\chi(z_{n})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}-\left(1-\frac{\chi(z_{k})_{+}}{\lvert\nabla\rho_{\varepsilon,k}\rvert}\right)_{+}\right|\leq\frac{\chi}{|\nabla\rho_{\varepsilon,k}|}|z_{n}-z_{k}|.$
Introducing this in (3.2) yields:
$\left(\int_{\mathbb{R}^{d}}\rvert\rho_{\varepsilon,n}-\rho_{\varepsilon,k}\lvert^{2}\,{\rm d}x\right)(t)\leq$ $\;\frac{\chi\sigma^{2}}{2\varepsilon}\int_{0}^{T}\int_{\mathbb{R}^{d}}\lvert(z_{n})_{+}-(z_{k})_{+}\rvert^{2}\,{\rm d}x\,{\rm d}t$
$-\frac{\varepsilon}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\lvert\nabla(\rho_{\varepsilon,n}-\rho_{\varepsilon,k})\rvert^{2}\,{\rm d}x\,{\rm d}t.$
We conclude that $\{\nabla\rho_{\varepsilon,n}\}$ is a Cauchy sequence in $L^{2}(0,T,L^{2}(\mathbb{R}^{d})))$, and therefore, $\nabla\rho_{\varepsilon,n}\rightarrow\nabla\rho_{\varepsilon}$ a.e. $x\in[0,T]\times\mathbb{R}^{d}$. The map $\mathcal{T}$ is continuous, compact and constant when $\sigma=0$. Moreover, fixed points are uniformly bounded in $\sigma$. Finally, we have to prove that fixed points of $\mathcal{T}$ are nonnegatives so we can remove the $+$ sign. Using Stampacchia’s truncation argument, we introduce $(\rho_{\varepsilon})_{-}$ in (11) as test function and we get
$\left(\int_{\mathbb{R}^{d}}(\rho_{\varepsilon})_{-}^{2}\,{\rm d}x\right)(t)\leq\int_{\mathbb{R}^{d}}(\rho_{{\rm in},\varepsilon})_{-}^{2}\,{\rm d}x.$
If $\rho_{{\rm in},\varepsilon}\geq 0$, solutions to $\mathcal{T}(\sigma,\rho_{\varepsilon})=\rho_{\varepsilon}$ are nonnegatives. Then, Leray-Schauder fixed-point yields a solution to
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\partial_{t}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t=$ $-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t$
$-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t,$ (21)
for $\forall\;\varphi\in L^{2}(0,T,H^{1}(\mathbb{R}^{d})$.
4. The limit $\varepsilon\rightarrow 0$ and proof of Theorem 1
The main difficulty in removing the viscosity terms $\varepsilon\Delta\rho_{\varepsilon}$ and $-\varepsilon\rho_{\varepsilon}$ is to pass to the limit in the nonlinear term $\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}\nabla\rho_{\varepsilon}$. We need to show that
$\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}\nabla\rho_{\varepsilon}\rightharpoonup\left(1-\frac{\chi\rho}{|\nabla\rho|}\right)_{+}\nabla\rho\quad\mbox{in }L^{2}(0,T;\mathbb{R}^{d}).$
We already know that
$\left\|\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}\nabla\rho_{\varepsilon}\right\|_{L^{2}(0,T;\mathbb{R}^{d})}\leq C,$
which implies,
$\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}\nabla\rho_{\varepsilon}\rightharpoonup G\quad\mbox{in }L^{2}(0,T;\mathbb{R}^{d}).$
The identification $G=\left(1-\frac{\chi\rho}{|\nabla\rho|}\right)_{+}\nabla\rho$ will follow by almost everywhere convergence of $\rho_{\varepsilon}$ (given by Lemma (4)) and of $\nabla\rho_{\varepsilon}$. The next two subsections will be devoted to proof that
$\nabla\rho_{\varepsilon}\rightarrow\nabla\rho\quad\mbox{a.e. in }\;{((0,T)\times\mathbb{R}^{d})\backslash\{|\nabla\rho|\leq\chi\rho\}.}$
Then, in the last two subsections, we will prove the $L^{p}$ convergence of $\rho$ to initial data $\rho_{{\rm in}}$ and the uniqueness of the solutions.
4.1. Entropy method
The major tool of this argument is the relative entropy between two functions $u$ and $v$, defined as:
$H_{{\rm rel}}[u|v]\equiv\int_{\mathbb{R}^{d}}\left(u\log\frac{u}{v}-u+v\right)\,{\rm d}x.$ (22)
Notice that, for $u,v\geq 0$
$H_{{\rm rel}}[u|v](T)\geq 0,\quad\mbox{as}\quad r\log(r)-r+1\geq 0,\quad\mbox{for all}\quad r\geq 0,$
and $H[u|v]=0$ if and only if $u=v$. Let us outline how we use (22) to show that
$\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}\nabla\rho_{\varepsilon}\rightharpoonup\left(1-\frac{\chi\rho}{|\nabla\rho|}\right)_{+}\nabla\rho\quad\mbox{in }L^{2}(0,T;\mathbb{R}^{d}).$
Consider $\rho_{\varepsilon}$ and $\rho_{\delta}$ two viscosity solutions to (3.2) with viscosity constants $\varepsilon$ and $\delta$ respectively. A formal computation shows that
$H[\rho_{\varepsilon}|\rho_{\delta}](T)-H[\rho_{{\rm in},\varepsilon}|\rho_{{\rm in},\delta}]+\int_{0}^{T}\mathcal{D}[\rho_{\varepsilon}|\rho_{\delta}]\,{\rm d}t=\mbox{error terms.}$. (23)
The relative entropy production $\mathcal{D}[\rho_{\varepsilon}|\rho_{\delta}]$ is nonnegative as a direct consequence of the monotonicity of the operator $\left(1-\frac{\chi}{|\nabla\log\rho|}\right)_{+}\nabla\log\rho$. The error terms originate from the viscosities $\varepsilon\Delta\rho_{\varepsilon}$ and $\delta\Delta\rho_{\delta}$ As $\varepsilon,\delta\rightarrow 0$, $H[\rho_{{\rm in},\varepsilon}|\rho_{{\rm in},\delta}]\rightarrow 0$ by construction, and the error terms also vanish. Since both $H$ and $\mathcal{D}$ are nonnegative, as $\delta,\varepsilon\to 0$, we must have
$H[\rho_{\varepsilon}|\rho_{\delta}](T)\rightarrow 0\quad\mbox{and}\quad\int_{0}^{T}\mathcal{D}[\rho_{\varepsilon}|\rho_{\delta}]\,{\rm d}t\rightarrow 0.$
From the convergence
$\int_{0}^{T}\mathcal{D}[\rho_{\varepsilon}|\rho_{\delta}]\,{\rm d}t\rightarrow 0,$
we obtain the almost everywhere convergence of $\nabla\rho_{\varepsilon}$ in the nondegenerate region $\{|\nabla\rho|>\chi\rho\}$. Consequently, we can identify the limit $\left(1-\frac{\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}\nabla\rho_{\varepsilon}$.
One can make this argument rigorous and obtain (23) by testing (3.2) for $\rho_{\varepsilon}$ and for $\rho_{\delta}$ respectively with test functions $\varphi_{1}=\log\left(\frac{\rho_{\varepsilon}}{\rho_{\delta}}\right)$ and $\varphi_{2}=\frac{\rho_{\varepsilon}}{\rho_{\delta}}$. Since neither $\varphi_{1}$ nor $\varphi_{2}$ are admissible test functions, we modify them accordingly, at the expense of some error terms. Specifically, our test functions are of the form
$\varphi_{1}=\log\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R},\;\;\mbox{and }\;\varphi_{2}=\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\phi_{R},$
with $\sigma>0$ a small constant and $\phi_{R}$ a cut-off function supported in $B_{R}$. More precisely, we chose $R=\sigma^{-\frac{1}{\alpha d}}$, with $\alpha>0$ to be fixed, and $\phi_{R}\in C^{\infty}_{c}(B_{2R})$ such that $\phi_{R}\equiv 1$ in $B_{R}$ and $|\nabla\phi_{R}|\leq\frac{C}{R}$ for some $C>0$. With this choice of $\varphi_{1}$ and $\varphi_{2}$ we obtain
$\int_{\mathbb{R}^{d}}$ $\left[\varphi_{1}\partial_{t}\rho_{\varepsilon}-\varphi_{2}\partial_{t}\rho_{\delta}\right]\,{\rm d}x=$
$=\frac{d}{dt}\int_{\mathbb{R}^{d}}\left[(\rho_{\varepsilon}+\sigma)\log{\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)}-\rho_{\varepsilon}+\rho_{\delta}\right]{{\phi_{R}}}\,{\rm d}x-\frac{d}{dt}\int_{\mathbb{R}^{d}}\rho_{\delta}\phi_{R}\,{\rm d}x.$ (24)
By Lemma 4, we know that
$\rho_{\varepsilon}\rightarrow\rho\quad\mbox{strongly in }L^{2}((0,T)\times\mathbb{R}^{d}).$
Hence, we can chose $\varepsilon,\delta>0$ small enough such that
$\int_{0}^{T}\int_{\mathbb{R}^{d}}|\rho_{\varepsilon}-\rho_{\delta}|^{2}\,{\rm d}x\,{\rm d}t<\sigma^{2}.$ (25)
We do not write $\varepsilon(\sigma),\delta(\sigma)$ and $R(\sigma)$ on $\sigma$ for readability purposes, but notice that
$\lim_{\sigma\rightarrow 0}\varepsilon,\delta=0,\quad\textrm{and}\quad\lim_{\sigma\rightarrow 0}R(\sigma)=\infty.$
Next, we subtract equation (3.2) for $\rho_{\delta}$, tested with $\varphi_{2}$, to the equation (3.2) for $\rho_{\varepsilon}$, tested with $\varphi_{1}$. This yields:
$\int_{0}^{T}$ $\int_{\mathbb{R}^{d}}\left[\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\partial_{t}\rho_{\varepsilon}-\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\;\partial_{t}\rho_{\delta}\right]\phi_{R}\,{\rm d}x\,{\rm d}t$ (26)
$=$ $-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\left[\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}+\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\nabla\phi_{R}\right]\,{\rm d}x\,{\rm d}t$
$+\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\nabla\rho_{\delta}\left[\nabla\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}+\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\nabla\phi_{R}\right]\,{\rm d}x\,{\rm d}t+E_{\varepsilon,\delta},$
where $E_{\varepsilon,\delta}$ are error terms, which must vanish when $\sigma\rightarrow 0$.
Note that (26) is an approximation of $\int_{0}^{T}\frac{d}{\,{\rm d}t}H_{{\rm rel}}[\rho_{\varepsilon}|\rho_{\delta}]$. We will show
$\lim\limits_{\varepsilon,\delta\rightarrow 0}\int_{0}^{T}\frac{d}{\,{\rm d}t}H_{{\rm rel}}[\rho_{\varepsilon}|\rho_{\delta}]\leq 0,$
to find that, in fact, $\lim\limits_{\varepsilon,\delta\rightarrow 0}H_{{\rm rel}}[\rho_{\varepsilon}|\rho_{\delta}](T)=0$. This limit will provide the necessary information about the convergence of $\nabla\rho_{\varepsilon}$.
To do this, let us analyze the error terms first:
$E_{\varepsilon,\delta}=$ $-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t+\delta\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\delta}\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t$ (27a)
$-\int_{0}^{T}\int_{\mathbb{R}^{d}}\phi_{R}\left[\varepsilon\nabla\rho_{\varepsilon}\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)-\delta\nabla\rho_{\delta}\nabla\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\right]\,{\rm d}x\,{\rm d}t$ (27b)
$-\int_{0}^{T}\int_{\mathbb{R}^{d}}\nabla\phi_{R}\left[\varepsilon\nabla\rho_{\varepsilon}\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)-\delta\nabla\rho_{\delta}\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\right]\,{\rm d}x\,{\rm d}t.$ (27c)
For the first term above, assuming without loss of generality $\varepsilon<\delta$ and recalling that we can always choose $\delta$ such that $\delta<\sigma<<1$ we compute
[…]
$+\delta\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\delta}\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t$
$\leq$ $\;-\varepsilon H[\rho_{\varepsilon}|\rho_{\delta}]+\sigma^{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t$ (28)
$+\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\delta}\phi_{R}\,{\rm d}x\,{\rm d}t+\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\phi_{R}\,{\rm d}x\,{\rm d}t$
$\leq$ $\;-\varepsilon H[\rho_{\varepsilon}|\rho_{\delta}]+\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(2\rho_{\varepsilon}+\rho_{\delta}+\phi_{R}\right)\,{\rm d}x\,{\rm d}t$
$\leq$ $\;-\varepsilon H[\rho_{\varepsilon}|\rho_{\delta}]+C\sigma^{\frac{\alpha-1}{\alpha}},$ (29)
where $C>0$ only depends on the $L^{1}$ norm of the initial data. In the last inequality we have used Lemma 1 and the fact that $\int_{\mathbb{R}^{d}}\phi_{R}\,{\rm d}x=C(d)\sigma^{\frac{\alpha-1}{\alpha}}$. From now on, $C$ will be a positive universal constant that can change from line to line.
For expression (27b), we compute
[…]
$-\delta\int_{0}^{T}\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\left|\frac{\nabla\rho_{\delta}}{\rho_{\delta}+\sigma}\right|^{2}\phi_{R}\,{\rm d}x\,{\rm d}t$
$\leq$ $\frac{(\varepsilon+\delta)}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\varepsilon}|^{2}}{\rho_{\varepsilon}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t+\frac{(\varepsilon+\delta)}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\left|\frac{\nabla\rho_{\delta}}{\rho_{\delta}+\sigma}\right|^{2}\phi_{R}\,{\rm d}x\,{\rm d}t$
$-\delta\int_{0}^{T}\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\left|\frac{\nabla\rho_{\delta}}{\rho_{\delta}+\sigma}\right|^{2}\phi_{R}\,{\rm d}x\,{\rm d}t$
$\leq$ $\frac{(\varepsilon+\delta)}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\varepsilon}|^{2}}{\rho_{\varepsilon}}\phi_{R}\,{\rm d}x\,{\rm d}t$
$\leq$ $\sigma C,$ (30)
using Lemma 3 in the last inequality. For the first term in (27c) we get
$\left|\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\nabla\phi_{R}\nabla\rho_{\varepsilon}\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\right|\,{\rm d}x\,{\rm d}t\leq$ $\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}|\log(\rho_{\varepsilon}+\sigma)||\nabla\phi_{R}||\nabla\rho_{\varepsilon}|\,{\rm d}x\,{\rm d}t$
$+\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}|\log(\rho_{\delta}+\sigma)||\nabla\phi_{R}||\nabla\rho_{\varepsilon}|\,{\rm d}x\,{\rm d}t.$
Since $|\log(\rho+\sigma)|\leq|\log(\sigma)|+(\rho+\sigma)$ and $|\nabla\phi_{R}|\leq C\sigma^{\frac{1}{\alpha d}}$, we estimate the two integrals above as
$\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}$ $\left|\log(\rho_{\varepsilon}+\sigma)\nabla\phi_{R}\nabla\rho_{\varepsilon}\right|\,{\rm d}x\,{\rm d}t$
$\leq$ $C\sigma^{\frac{1}{\alpha d}+1}\int_{0}^{T}\int_{B_{2R}}\left(|\log(\sigma)||\nabla\rho_{\varepsilon}|+|\nabla\rho_{\varepsilon}|(\rho_{\varepsilon}+\sigma)\right)\,{\rm d}x\,{\rm d}t$
$\leq$ $\;C\sigma^{\frac{1}{\alpha d}+1}|\log(\sigma)|\int_{0}^{T}\int_{B_{2R}}\left(\frac{|\nabla\rho_{\varepsilon}|^{2}}{\rho_{\varepsilon}}+\rho_{\varepsilon}\right)\,{\rm d}x\,{\rm d}t$
$+\frac{C}{2}\sigma^{\frac{1}{\alpha d}+1}\int_{0}^{T}\int_{B_{2R}}\left(\sigma\frac{|\nabla\rho_{\varepsilon}|^{2}}{\rho_{\varepsilon}}+\sigma\rho_{\varepsilon}+|\nabla\rho_{\varepsilon}|^{2}+\rho_{\varepsilon}^{2}\right)$
$\leq$ $\;C\sigma,$ (31)
for sufficiently small $\sigma$, by Lemma 1 and Proposition 3 in the last inequality. Similar computations can be done for the integral involving $\rho_{\delta}$.
For the second term in (27c), we compute
$\left|\delta\int_{0}^{T}\int_{\mathbb{R}^{d}}\nabla\phi_{R}\nabla\rho_{\delta}\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\,{\rm d}x\,{\rm d}t\right|\leq$ $\;\sigma^{\frac{1}{\alpha d}}\int_{0}^{T}\int_{\mathbb{R}^{d}}|\nabla\rho_{\delta}|(\rho_{\varepsilon}+\sigma)\,{\rm d}x\,{\rm d}t$
$\leq$ $\;\;\frac{\sigma^{\frac{1}{\alpha d}}}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(|\nabla\rho_{\delta}|^{2}+\rho_{\varepsilon}^{2}+\sigma\frac{|\nabla\rho_{\delta}|^{2}}{\rho_{\delta}}+\sigma\rho_{\delta}\right)\,{\rm d}x\,{\rm d}t$
$\leq$ $\;\sigma^{\frac{1}{\alpha d}}C,$ (32)
using Proposition 1 in the last inequality. Summarizing, by (4.1), (4.1), (4.1) and (4.1), for a sufficiently small $\sigma$, we have the following estimate for the error term
$E_{\varepsilon,\delta}\leq-\varepsilon H[\rho_{\varepsilon}|\rho_{\delta}]+C\,\sigma^{{\frac{1}{\alpha d}}}.$ (33)
Now, we have to estimate the following two terms from (26) involving $\nabla\phi_{R}$:
${I_{1}:=}$ $-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\nabla\phi_{R}\log\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\,{\rm d}x\,{\rm d}t,$ (34a)
${I_{2}:=}$ $\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\nabla\rho_{\delta}\nabla\phi_{R}\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\,{\rm d}x\,{\rm d}t.$ (34b)
Following the same strategy as in (4.1), for (34a) we get that
$\left|{I_{1}}\right|\lesssim\;C\sigma^{\frac{1}{\alpha d}}|\log(\sigma)|.$
Similarly, (34b) can be estimated as
$\left|{I_{2}}\right|$ $\leq\sigma^{\frac{1}{\alpha d}}\int_{0}^{T}\int_{\mathbb{R}^{d}}|\nabla\rho_{\delta}|\left(1+\frac{|\rho_{\varepsilon}-\rho_{\delta}|}{\rho_{\delta}+\sigma}\right)\,{\rm d}x\,{\rm d}t$
$\leq\sigma^{\frac{1}{\alpha d}}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(\frac{|\nabla\rho_{\delta}|^{2}}{\rho_{\delta}}+\rho_{\delta}+|\nabla\rho_{\delta}|^{2}+\frac{1}{\sigma^{2}}|\rho_{\varepsilon}-\rho_{\delta}|^{2}\right)\,{\rm d}x\,{\rm d}t$
$\leq\sigma^{\frac{1}{\alpha d}}C,$
by a priori estimates and (25).
Now, let us see what happens with the remaining term in (26). These are
${I_{3}}:=-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}\,{\rm d}x\,{\rm d}t$ (35)
${I_{4}}:=\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\nabla\rho_{\delta}\nabla\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}\,{\rm d}x\,{\rm d}t.$ (36)
We can reformulate $I_{3}$ as
${I_{3}}=$ $-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}\,{\rm d}x\,{\rm d}t$
$-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\right]\nabla\rho_{\varepsilon}\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}\,{\rm d}x\,{\rm d}t.$
We rewrite $I_{4}$ in a similar way. It follows
${I_{3}+I_{4}}=$ $\;-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}\,{\rm d}x\,{\rm d}t$
$+\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\nabla\rho_{\delta}\nabla\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}\,{\rm d}x\,{\rm d}t+R_{1}+R_{2},$
where
$R_{1}:=-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\right]\nabla\rho_{\varepsilon}\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}\,{\rm d}x\,{\rm d}t,$
$R_{2}:=\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\nabla\rho_{\delta}\nabla\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}\,{\rm d}x\,{\rm d}t.$
We recall that
$\left|\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\right|\leq\frac{\chi\sigma}{|\nabla\rho_{\varepsilon}|},$ (37)
and also that
$0\leq\left[\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\right].$
Similar bounds hold in $R_{2}$ for the term with $\rho_{\delta}$. Using this, we estimate $R_{1}$ as follows:
$R_{1}\leq$ $\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\right]\nabla\rho_{\varepsilon}\frac{\nabla\rho_{\delta}}{\rho_{\delta}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t$
$\leq$ $\chi\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\delta}|}{\rho_{\delta}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t\leq\chi\sigma^{1/2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\delta}|}{\sqrt{\rho_{\delta}+\sigma}}\phi_{R}\,{\rm d}x\,{\rm d}t$
$\leq$ $\chi\sigma^{\frac{1}{2}-\frac{1}{2\alpha}}C,$
which vanishes when $\sigma\rightarrow 0$ if $\alpha>1$. We estimate $R_{2}$ in a similar way:
$R_{2}\leq$ $\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\frac{\nabla\rho_{\delta}\nabla\rho_{\varepsilon}}{\rho_{\delta}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t$
$\leq$ $\chi\sigma\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\varepsilon}|}{\rho_{\delta}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t\leq\chi\sigma^{1/2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\varepsilon}|\sqrt{\rho_{\varepsilon}}}{\sqrt{\rho_{\varepsilon}}\sqrt{\rho_{\delta}+\sigma}}\phi_{R}\,{\rm d}x\,{\rm d}t$
$\leq$ $\chi\sigma^{1/2}\left(\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\varepsilon}|}{\rho_{\varepsilon}}\phi_{R}\,{\rm d}x\,{\rm d}t\right)^{1/2}\left(\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}}{\rho_{\delta}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t\right)^{1/2}$
$\leq$ $C\sigma^{1/2}\left(\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\rho_{\varepsilon}-\rho_{\delta}|}{\rho_{\varepsilon}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t+\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\rho_{\delta}}{\rho_{\delta}+\sigma}\phi_{R}\,{\rm d}x\,{\rm d}t\right)^{1/2}$
$\leq$ $C\sigma^{1/2}\left(\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\rho_{\varepsilon}-\rho_{\delta}|^{2}}{\sigma^{2}}\,{\rm d}x\,{\rm d}t+\int_{0}^{T}\int_{\mathbb{R}^{d}}\phi_{R}^{2}\,{\rm d}x\,{\rm d}t+\int_{0}^{T}\int_{\mathbb{R}^{d}}\phi_{R}\,{\rm d}x\,{\rm d}t\right)^{1/2}$
$\leq$ $C\sigma^{1/2}\left(1+\int_{0}^{T}\int_{\mathbb{R}^{d}}(\phi_{R}+\phi_{R}^{2})\,{\rm d}x\,{\rm d}t\right)^{1/2}$
$\leq$ $C\sigma^{\frac{1}{2}-\frac{1}{2\alpha}}.$
We conclude that, if $\alpha>1$, $R_{1},R_{2}\rightarrow 0$ when $\sigma\rightarrow 0$.
Straightforward computations lead to
${I_{3}+I_{4}}=$ $-\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}+\sigma}{2}\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\times$
×$\times\left(\lvert\nabla\log(\rho_{\varepsilon}+\sigma)\rvert^{2}-\lvert\nabla\log(\rho_{\delta}+\sigma)\rvert^{2}\right)\,{\rm d}x\,{\rm d}t$
$-\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}+\sigma}{2}\left|\nabla\log\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\right|^{2}\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}+\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\,{\rm d}x\,{\rm d}t$
${+R_{1}+R_{2}}$
$\leq$ $\;-\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}+\sigma}{2}\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{|\nabla\rho_{\varepsilon}|}\right)_{+}-\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{|\nabla\rho_{\delta}|}\right)_{+}\right]^{2}\,{\rm d}x\,{\rm d}t$
$-\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\rho_{\varepsilon}+\sigma}{2}\left|\nabla\log\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\right|^{2}\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}+\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\,{\rm d}x\,{\rm d}t$
$+R_{1}+R_{2},$
where, in the second inequality, we have used that
$\left[\left(1-\frac{\chi}{r}\right)_{+}-\left(1-\frac{\chi}{s}\right)_{+}\right]^{2}\leq C\left[\left(1-\frac{\chi}{r}\right)_{+}-\left(1-\frac{\chi}{s}\right)_{+}\right]\left(r^{2}-s^{2}\right),\quad\mbox{for any}\quad r,s>0.$
Therefore, (4.1) and (26) give
$\left(\int_{\mathbb{R}^{d}}\left[(\rho_{\varepsilon}+\sigma)\log{\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)}-\rho_{\varepsilon}+\rho_{\delta}\right]\,{\rm d}x\right)(T)$
$\qquad{-\int_{\mathbb{R}^{d}}\left[(\rho_{in,\varepsilon}+\sigma)\log{\left(\frac{\rho_{in,\varepsilon}+\sigma}{\rho_{in,\delta}+\sigma}\right)}-\rho_{in,\varepsilon}+\rho_{in,\delta}\right]\,{\rm d}x}$
$={\int_{\mathbb{R}^{d}}(\rho_{\delta}(T)-\rho_{in,\delta})\phi_{R}\,{\rm d}x}$
$\;\;\;-\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{|\nabla\rho_{\varepsilon}|}\right)_{+}-\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{|\nabla\rho_{\delta}|}\right)_{+}\right]^{2}\,{\rm d}x\,{\rm d}t$
$\;\;\;-\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\left|\nabla\log\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\right|^{2}\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}+\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\,{\rm d}x\,{\rm d}t$
$\;\;\;+R_{1}+R_{2}+{I_{1}+I_{2}}+E_{\varepsilon,\delta},$ (38)
where, $R_{1}$, $R_{2}$, $I_{1}$, $I_{2}$ and $E_{\varepsilon,\delta}$ vanish when $\sigma,\varepsilon,\delta\rightarrow 0$. Now, we need to prove
${\int_{\mathbb{R}^{d}}(\rho_{\delta}(T)-\rho_{in,\delta})\phi_{R}\,{\rm d}x}\rightarrow 0$ when $\sigma,\varepsilon,\delta\rightarrow 0$. For this, we test (3.2) with $\phi_{R}$ to get
$\left(\int_{\mathbb{R}^{d}}\rho_{\delta}{\phi_{R}}\,{\rm d}x\right)(T)=$ $\int_{\mathbb{R}^{d}}\rho_{{\rm in},\delta}{\phi_{R}}\,{\rm d}x-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\nabla\rho_{\delta}\nabla\phi_{R}\,{\rm d}x\,{\rm d}t$
$-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\nabla\rho_{\delta}\nabla\phi_{R}\,{\rm d}x\,{\rm d}t-\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\delta}\phi_{R}\,{\rm d}x\,{\rm d}t.$ (39)
Since $|\nabla\phi_{R}|\leq\frac{C}{R(\sigma)}\rightarrow 0$ and $\delta\rightarrow 0$ when $\sigma\rightarrow 0$, we have that
$\lim_{\sigma\rightarrow 0}\left\lvert\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\delta}}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\nabla\rho_{\delta}\nabla\phi_{R}\,{\rm d}x\,{\rm d}t\right\rvert\leq\lim_{\sigma\rightarrow 0}\frac{C}{R(\sigma)}\|\nabla\sqrt{\rho_{\delta}}\|_{2}\|\rho_{\delta}\|_{1}=0.$
The viscosity terms in (4.1) also vanish when $\sigma\rightarrow 0$, and therefore we get that
$\lim_{\sigma\rightarrow 0}\left(\int_{\mathbb{R}^{d}}\rho_{\delta}\varphi\,{\rm d}x\right)(T)=\lim_{\sigma\rightarrow 0}\int_{\mathbb{R}^{d}}\rho_{{\rm in},\delta}\varphi\,{\rm d}x.$
Finally, taking the limit $\sigma,\varepsilon,\delta\rightarrow 0$ in (4.1) yields
$\lim_{\epsilon,\delta\rightarrow 0}\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{|\nabla\rho_{\varepsilon}|}\right)_{+}-\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{|\nabla\rho_{\delta}|}\right)_{+}\right]^{2}\phi_{R}\,{\rm d}x\,{\rm d}t=0,$ (40a)
$\lim_{\epsilon,\delta\rightarrow 0}\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}(\rho_{\varepsilon}+\sigma)\left|\nabla\log\left(\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\right|^{2}\cdot$
$\;\;\;\;\;\;\;\;\;\;\;\;\cdot\left[\left(1-\frac{\chi(\rho_{\varepsilon}+\sigma)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}+\left(1-\frac{\chi(\rho_{\delta}+\sigma)}{\lvert\nabla\rho_{\delta}\rvert}\right)_{+}\right]\phi_{R}\,{\rm d}x\,{\rm d}t=0.$ (40b)
4.2. Identification of the limit of $\nabla\rho_{\varepsilon}$
Now, from (40a) is easy to see that
$\left(1-\frac{\chi(\rho_{\varepsilon}+\varepsilon)}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\rightarrow\overline{G}\quad\mbox{a.e. }x\in[0,T]\times\mathbb{R}^{d}\setminus\{\rho=0\}.$
And on the other hand, from (40b)
$\frac{\nabla\rho_{\varepsilon}}{\rho_{\varepsilon}+\varepsilon}\rightarrow w\quad\mbox{a.e. in }[0,T]\times\mathbb{R}^{d}\setminus\{\{\rho=0\}\cup\{\overline{G}=0\}\}.$
Notice, moreover, that if $\left(1-\frac{\chi(\rho_{\varepsilon}+\varepsilon)}{|\nabla\rho_{\varepsilon}|}\right)_{+}$ converges almost everywhere in $[0,T]\times\mathbb{R}^{d}\setminus\{\rho=0\}$ then
$\frac{\left|\nabla\rho_{\varepsilon}\right|}{\rho_{\varepsilon}+\varepsilon}<\infty\quad\mbox{a.e. in }[0,T]\times\mathbb{R}^{d}\setminus\{\rho=0\},$
and we have
$(\rho_{\varepsilon}+\varepsilon)^{3}\frac{\left|\nabla\rho_{\varepsilon}\right|}{\rho_{\varepsilon}+\varepsilon}\left(\frac{\left|\nabla\rho_{\varepsilon}\right|}{\rho_{\varepsilon}+\varepsilon}+\chi\right)\left(1-\frac{\chi(\rho_{\varepsilon}+\varepsilon)}{|\nabla\rho_{\varepsilon}|}\right)\frac{\nabla\rho_{\varepsilon}}{\rho_{\varepsilon}+\varepsilon}\rightarrow\overline{w},$
almost everywhere in $[0,T]\times\mathbb{R}^{d}\setminus\{\rho=0\}$, or, which is equivalent,
$\left(|\nabla\rho_{\varepsilon}|^{2}-(\rho_{\varepsilon}+\varepsilon)^{2}\chi^{2}\right)_{+}\nabla\rho_{\varepsilon}\rightarrow\overline{w},$
almost everywhere in $[0,T]\times\mathbb{R}^{d}\setminus\{\rho=0\}$. But as $\rho_{\varepsilon}\rightarrow\rho$ strongly in $L^{2}(0,T,\mathbb{R}^{d})$, $\rho_{\varepsilon}\rightarrow\rho$ a.e. in $(0,T)\times\mathbb{R}^{d}$, and
$\left(|\nabla\rho_{\varepsilon}|^{2}-\rho^{2}\chi^{2}\right)_{+}\nabla\rho_{\varepsilon}\rightarrow\overline{w},$
almost everywhere in $[0,T]\times\mathbb{R}^{d}\setminus\{\rho=0\}$.
Define $Q_{R,\tau,T}=\{[\tau,T]\times B_{R}\cap\{\rho>0\}\}$. Egorov’s and Lusin’s Theorems imply that, for every $\eta>0$, there exists $Q_{R,\tau,T}^{\eta}\subset Q_{R,\tau,T}$ compact such that $|Q_{R,\tau,T}\backslash Q_{R,\tau,T}^{\eta}|<\eta$,
$\left(|\nabla\rho_{\varepsilon}|^{2}-\rho^{2}\chi^{2}\right)_{+}\nabla\rho_{\varepsilon}\in C^{0}(Q_{R,\tau,T}^{\eta}),$
$\left(|\nabla\rho_{\varepsilon}|^{2}-\rho^{2}\chi^{2}\right)_{+}\nabla\rho_{\varepsilon}\to\overline{w}\quad\mbox{uniformly in }Q_{R,\tau,T}^{\eta}.$
Then, as $\nabla\rho_{\varepsilon}\rightharpoonup\nabla\rho$ weakly in $L^{p}(\mathbb{R}^{d}\times(0,T))$ we have that
$\int_{Q_{R,\tau,T}^{\eta}}\left[\left(|\nabla\rho_{\varepsilon}|^{2}-\rho^{2}\chi^{2}\right)_{+}\nabla\rho_{\varepsilon}-\left(|\nabla\rho|^{2}-\rho^{2}\chi^{2}\right)_{+}\nabla\rho\right](\nabla\rho_{\varepsilon}-\nabla\rho)\,{\rm d}x\,{\rm d}t$
$=$ $\int_{Q_{R,\tau,T}^{\eta}}\left[\left(|\nabla\rho_{\varepsilon}|^{2}-\rho^{2}\chi^{2}\right)_{+}\nabla\rho_{\varepsilon}-\overline{w}\right](\nabla\rho_{\varepsilon}-\nabla\rho)\,{\rm d}x\,{\rm d}t$
$+\int_{Q_{R,\tau,T}^{\eta}}\left[\overline{w}-\left(|\nabla\rho|^{2}-\rho^{2}\chi^{2}\right)_{+}\nabla\rho\right](\nabla\rho_{\varepsilon}-\nabla\rho)\,{\rm d}x\,{\rm d}t\rightarrow 0.$
But, notice that
$\int_{Q_{R,\tau,T}^{\eta}}\left[\left(|\nabla\rho_{\varepsilon}|^{2}-\rho^{2}\chi^{2}\right)_{+}\nabla\rho_{\varepsilon}-\left(|\nabla\rho|^{2}-\rho^{2}\chi^{2}\right)_{+}\nabla\rho\right](\nabla\rho_{\varepsilon}-\nabla\rho_{\varepsilon})\,{\rm d}x\,{\rm d}t$
$=$ $\frac{1}{2}\int_{Q_{R,\tau,T}^{\eta}}\left[\left(|\nabla\rho_{\varepsilon}|^{2}-\rho^{2}\chi^{2}\right)_{+}+\left(|\nabla\rho|^{2}-\rho^{2}\chi^{2}\right)_{+}\right]\times\lvert\nabla(\rho_{\varepsilon}-\rho)\rvert^{2}\,{\rm d}x\,{\rm d}t$ (41a)
$+\frac{1}{2}\int_{Q_{R,\tau,T}^{\eta}}\left[\left(|\nabla\rho_{\varepsilon}|^{2}-\rho^{2}\chi^{2}\right)_{+}-\left(|\nabla\rho|^{2}-\rho^{2}\chi^{2}\right)_{+}\right]\times\left(\lvert\nabla\rho_{\varepsilon}\rvert^{2}-\lvert\nabla\rho_{\varepsilon}\rvert^{2}\right)\,{\rm d}x\,{\rm d}t\rightarrow 0.$ (41b)
Since both integrals in (41a), (41b) are nonnegative, the convergence to 00 of (41a) implies that we can find a subsequence (not relabeled) such that $\nabla\rho_{\varepsilon}\rightarrow\nabla\rho$ a.e. in $Q_{R,\tau,T}^{\eta}\setminus\{\rho\chi\geq|\nabla\rho|\}$. W.l.o.g. we can choose $\eta=1/N$, $N\in\mathbb{N}$, and assume that $Q_{R,\tau,T}^{1/N_{1}}\subset Q_{R,\tau,T}^{1/N_{2}}$ for $N_{1}<N_{2}$. Since $|Q_{R,\tau,T}\setminus Q_{R,\tau,T}^{\eta}|<\eta$, a Cantor diagonal argument yields the existence of a subsequence (not relabeled) such that $\nabla\rho_{\varepsilon}\rightarrow\nabla\rho$ a.e. in $Q_{R,\tau,T}\setminus\{\rho\chi\geq|\nabla\rho|\}$. Via another Cantor diagonal argument we conclude that (up to a subsequence)
$\nabla\rho_{\varepsilon}\rightarrow\nabla\rho\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{ a.e. in }(0,T)\times\mathbb{R}^{d}\setminus\{\rho\chi\geq|\nabla\rho|\}.$ (42)
Thus
$\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\rightharpoonup\left(1-\frac{\chi\rho}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho\quad\mbox{weakly in }L^{2}((0,T)\times\mathbb{R}^{d}\setminus\{\rho\chi\geq|\nabla\rho|\}).$
Furthermore
$\left[\left(|\nabla\rho_{\varepsilon}|^{2}-\rho^{2}\chi^{2}\right)_{+}-\left(|\nabla\rho|^{2}-\rho^{2}\chi^{2}\right)_{+}\right]\times\left(\lvert\nabla\rho_{\varepsilon}\rvert^{2}-\lvert\nabla\rho_{\varepsilon}\rvert^{2}\right)$
$=$ $\left[\left(|\nabla\rho_{\varepsilon}|^{2}-\rho^{2}\chi^{2}\right)_{+}-\left(|\nabla\rho|^{2}-\rho^{2}\chi^{2}\right)_{+}\right]^{2}$
$-\left[\left(|\nabla\rho_{\varepsilon}|^{2}-\rho^{2}\chi^{2}\right)_{+}-\left(|\nabla\rho|^{2}-\rho^{2}\chi^{2}\right)_{+}\right]\left[\left(|\nabla\rho_{\varepsilon}|^{2}-\rho^{2}\chi^{2}\right)_{-}-\left(|\nabla\rho|^{2}-\rho^{2}\chi^{2}\right)_{-}\right]$
$=$ $\left[\left(|\nabla\rho_{\varepsilon}|^{2}-\rho^{2}\chi^{2}\right)_{+}-\left(|\nabla\rho|^{2}-\rho^{2}\chi^{2}\right)_{+}\right]^{2}$
$+\left(|\nabla\rho_{\varepsilon}|^{2}-\rho^{2}\chi^{2}\right)_{+}\left(|\nabla\rho|^{2}-\rho^{2}\chi^{2}\right)_{-}+\left(|\nabla\rho_{\varepsilon}|^{2}-\rho^{2}\chi^{2}\right)_{-}\left(|\nabla\rho|^{2}-\rho^{2}\chi^{2}\right)_{+}.$
So, (41b) gives
$\left(|\nabla\rho_{\varepsilon}|^{2}-\rho_{\varepsilon}^{2}\chi^{2}\right)_{+}\rightarrow\left(|\nabla\rho|^{2}-\rho^{2}\chi^{2}\right)_{+}\quad\mbox{a.e. in }L^{2}({(0,T)\times\mathbb{R}^{d}\cap\{\rho>0\}}).$
Furthermore, notice that
$\lim\limits_{\varepsilon\rightarrow 0}\left|\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\right|\left(|\nabla\rho_{\varepsilon}|+\rho_{\varepsilon}\chi\right)_{+}\leq\lim\limits_{\varepsilon\rightarrow 0}\left(|\nabla\rho_{\varepsilon}|^{2}-\rho_{\varepsilon}^{2}\chi^{2}\right)_{+}=\left(|\nabla\rho|^{2}-\rho^{2}\chi^{2}\right)_{+}=0,$
a.e. in $(0,T)\times\mathbb{R}^{d}\cap\{\rho\chi\geq|\nabla\rho,|\leavevmode\nobreak\ {\rho>0}\}$. Finally, it holds
$\left|\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\right|\leq 2\sqrt{\rho_{\varepsilon}}|\nabla\sqrt{\rho_{\varepsilon}}|\to 0=\left|\left(1-\frac{\chi\rho}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho\right|\quad\mbox{a.e.\leavevmode\nobreak\ in }\{\rho>0\}.$
This means that
$\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\rightarrow\left(1-\frac{\chi\rho}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho\quad\mbox{a.e. in }(0,T)\times\mathbb{R}^{d},$
and we finally get
$\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\rightharpoonup\left(1-\frac{\chi\rho}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho\quad\mbox{weakly in }L^{2}((0,T)\times\mathbb{R}^{d}).$
4.3. Limit $\varepsilon\rightarrow 0$ in (3.2)
As in (19), we have that
$\|\partial\rho_{\varepsilon}\|_{L^{2}(0,T;H^{-1}(\mathbb{R}^{d}))}\leq C\|\rho_{{\rm in},\varepsilon}\|_{L^{2}(\mathbb{R}^{d})},$
and, as $\rho_{{\rm in},\varepsilon}$ is uniformly bounded in $L^{2}(0,T;\mathbb{R}^{d})$, this yields
$\partial_{t}\rho_{\varepsilon}\rightharpoonup\partial_{t}\rho\quad\mbox{weakly in }\;\;L^{2}(0,T;H^{-1}).$
On the other hand, due to what is stated in Section 4, we have that
$\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\rightharpoonup\left(1-\frac{\chi\rho}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho\quad\mbox{weakly in }\;\;L^{2}(0,T;\mathbb{R}^{d}).$
For the regularizing terms we have the following:
$\lim_{\varepsilon\rightarrow 0}\left|\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\nabla\rho_{\varepsilon}\nabla\varphi\,{\rm d}x\,{\rm d}t+\varepsilon\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t\right|$ $\leq\lim_{\varepsilon\rightarrow 0}\varepsilon C\|\rho_{{\rm in},\varepsilon}\|_{L^{2}(0,T;\mathbb{R}^{d})}\|\varphi\|_{L^{2}(0,T;\mathbb{R}^{d})}$
$=0.$
Finally we have that $\rho$ is the weak solution to (10).
4.4. Obtaining the initial data $\rho_{{\rm in}}$ for $\rho$
In what follows, we will show that $\rho(t)\rightarrow\rho_{{\rm in}}$ strongly in $L^{p}(\mathbb{R}^{d})$ as $t\to 0$.
We start by reminding that $\rho_{\varepsilon}$ is bounded in the space
$\left\{u\in L^{2}(0,T;H^{1}(\mathbb{R}^{d}))\leavevmode\nobreak\ \leavevmode\nobreak\ :\leavevmode\nobreak\ \leavevmode\nobreak\ \partial_{t}u\in L^{2}(0,T;H^{-1}(\mathbb{R}^{d}))\right\}$
which embeds continuously into $C^{0}([0,T];L^{2}(\mathbb{R}^{d}))$ [17, Prop. 23.23]. In particular it follows that $\rho_{\varepsilon}(t)\rightharpoonup\rho(t)$ weakly in $H^{-1}(\mathbb{R}^{d})$ for every $t\in[0,T]$, hence
[…]
Also remember that the initial condition for the approximated problem (3.2) is satisfied in the sense of (18).
Define for brevity $A_{\varepsilon}:=\varepsilon+\left(1-\frac{\chi\rho_{\varepsilon}}{|\nabla\rho_{\varepsilon}|}\right)_{+}$, $A:=\left(1-\frac{\chi\rho}{|\nabla\rho|}\right)_{+}$. Choose $\varphi\in H^{1}(\mathbb{R}^{d})$ independent of time in (3.2). It follows
$\left|\int_{\mathbb{R}^{d}}\rho_{\varepsilon}(\tau)\varphi\,{\rm d}x-\int_{\mathbb{R}^{d}}\rho_{{\rm in},\varepsilon}\varphi\,{\rm d}x\right|$
$\quad=\left|\int_{0}^{\tau}\int_{\mathbb{R}^{d}}A_{\varepsilon}\nabla\rho_{\varepsilon}\cdot\nabla\varphi\,{\rm d}x\,{\rm d}t+\varepsilon\int_{0}^{\tau}\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\varphi\,{\rm d}x\,{\rm d}t\right|$
$\quad\leq(1+\varepsilon)\left(\|\nabla\rho_{\varepsilon}\|_{L^{2}(0,\tau;L^{2}(\mathbb{R}^{d}))}\|\nabla\varphi\|_{L^{2}(0,\tau;L^{2}(\mathbb{R}^{d}))}+\|\rho_{\varepsilon}\|_{L^{2}(0,\tau;L^{2}(\mathbb{R}^{d}))}\|\varphi\|_{L^{2}(0,\tau;L^{2}(\mathbb{R}^{d}))}\right)$
$\quad\leq C\tau^{1/2}\|\varphi\|_{H^{1}(\mathbb{R}^{d})}.$
Relation (43) allows us to take the limit $\varepsilon\to 0$ in the above estimate and obtain
$\left|\int_{\mathbb{R}^{d}}\rho(\tau)\varphi\,{\rm d}x-\int_{\mathbb{R}^{d}}\rho_{0}\varphi\,{\rm d}x\right|\leq C\tau^{1/2}\|\varphi\|_{H^{1}(\mathbb{R}^{d})},$
implying
$\rho(t)\to\rho_{0}\quad\mbox{strongly in $H^{-1}(\mathbb{R}^{d})$ as $t\to 0$.}$ (44)
Define now, for $M>0$ arbitrary,
[…]
Choose $\varphi=f_{M}(\rho){\raisebox{3.0pt}{\Large$\chi$}}_{[t_{1},t_{2}]}(t)\in L^{2}(0,T;H^{1}(\mathbb{R}^{d}))$ in (10). It follows
[…]
Since $f_{M}^{\prime}(\rho)\rho^{2-p}=p(p-1)\min(\rho^{p-2},M^{p-2})\rho^{2-p}\leq p(p-1)$, $F_{M}(\rho)\leq\rho^{p}$, Proposition 1 and the dominated convergence theorem allow us to take the limit $M\to\infty$ in the above identity and deduce
[…]
for a.e. $t_{1},t_{2}\in[0,T]$. It follows that the function $t\mapsto\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}^{p}$ is in $W^{1,1}(0,T)$, hence up to changing the values of $\rho:[0,T]\to L^{p}(\mathbb{R}^{d})$ in a zero measure set it holds that $t\mapsto\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}^{p}$ is continuous in $[0,T]$. As a consequence the bound $\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}\leq\|\rho_{0}\|_{L^{p}(\mathbb{R}^{d})}$ holds for every $t\in[0,T]$. In particular $\rho(t)$ is weakly convergent (up to subsequences) in $L^{p}(\mathbb{R}^{d})$ as $t\to 0$; however, its weak limit has to be $\rho_{0}$ due to (44), hence $\rho(t)\rightharpoonup\rho_{0}$ weakly in $L^{p}(\mathbb{R}^{d})$ as $t\to 0$ (no longer up to subsequences). It follows
$\|\rho_{0}\|_{L^{p}(\mathbb{R}^{d})}\leq\liminf_{t\to 0}\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}\leq\limsup_{t\to 0}\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}\leq\|\rho_{0}\|_{L^{p}(\mathbb{R}^{d})}$
implying that $\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}\to\|\rho_{0}\|_{L^{p}(\mathbb{R}^{d})}$ as $t\to 0$. This fact and the weak convergence $\rho(t)\rightharpoonup\rho_{0}$ in $L^{p}(\mathbb{R}^{d})$, imply that $\rho(t)\rightarrow\rho_{{\rm in}}$ strongly in $L^{p}(\mathbb{R}^{d})$ as $t\to 0$.
4.5. Uniqueness
Let $\rho_{1}$ and $\rho_{2}$ be two solutions to (10) with the same initial data. As in Section 4, we subtract equation (10) for $\rho_{2}$ tested with $\varphi_{2}=\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}{-1}=\frac{\rho_{1}-\rho_{2}}{\rho_{2}+\sigma}$ to the equation for $\rho_{1}$ tested with $\varphi_{1}=\log\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)$. Following the same computations as in (4.1), we get
$\int_{0}^{T}\int_{\mathbb{R}^{d}}$ $\left[\varphi_{1}\partial_{t}\rho_{1}-\varphi_{2}\partial_{t}\rho_{2}\right]\,{\rm d}x\;dt=$
$=\int_{0}^{T}\frac{d}{dt}\int_{\mathbb{R}^{d}}(\rho_{2}+\sigma)\left[\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)\log{\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)}-\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)+1\right]\,{\rm d}x\,{\rm d}t$
$=\int_{\mathbb{R}^{d}}(\rho_{2}+\sigma)\left[\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)\log{\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)}-\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)+1\right](T)\,{\rm d}x$
After some manipulations, this yields
$\int_{\mathbb{R}^{d}}(\rho_{2}+\sigma)\left[\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)\log{\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)}-\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)+1\right](T)\,{\rm d}x$ (45)
$=-\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}(\rho_{1}+\sigma)\left[\left(1-\frac{\chi(\rho_{1}+\sigma)}{\lvert\nabla\rho_{1}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\cdot$
$\;\;\;\;\;\;\;\;\;\;\cdot\left(\lvert\nabla\log(\rho_{1}+\sigma)\rvert^{2}-\lvert\nabla\log(\rho_{2}+\sigma)\rvert^{2}\right)\,{\rm d}x\,{\rm d}t$
$\;\;\;-\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}(\rho_{1}+\sigma)\left|\nabla\log\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)\right|^{2}\left[\left(1-\frac{\chi(\rho_{1}+\sigma)}{\lvert\nabla\rho_{1}\rvert}\right)_{+}+\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\,{\rm d}x\,{\rm d}t$
$\;\;\;+\xi_{1}+\xi_{2}$
${\leq\xi_{1}+\xi_{2}}$
where
$\xi_{1}:=-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{1}}{\lvert\nabla\rho_{1}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{1}+\sigma)}{\lvert\nabla\rho_{1}\rvert}\right)_{+}\right]\nabla\rho_{1}\nabla\left(\log\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)\,{\rm d}x\,{\rm d}t,$
and
$\xi_{2}:=\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{2}}{\lvert\nabla\rho_{2}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\nabla\rho_{2}\nabla\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)\,{\rm d}x\,{\rm d}t.$
We want to show that $\limsup_{\sigma\to 0}(\xi_{1}+\xi_{2})\leq 0$. We estimate first $\xi_{1}$ and then $\xi_{2}$.
$\xi_{1}\leq$ $\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{1}}{\lvert\nabla\rho_{1}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{1}+\sigma)}{\lvert\nabla\rho_{1}\rvert}\right)_{+}\right]\nabla\rho_{1}\frac{\nabla\rho_{2}}{\rho_{2}+\sigma}\,{\rm d}x\,{\rm d}t$
$\leq$ $\underbrace{\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{1}}{\lvert\nabla\rho_{1}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{1}+\sigma)}{\lvert\nabla\rho_{1}\rvert}\right)_{+}\right]\frac{|\nabla\rho_{1}|^{2}}{\rho_{2}+\sigma}\,{\rm d}x\,{\rm d}t}_{\text{(I)}}$
$+\underbrace{\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{1}}{\lvert\nabla\rho_{1}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{1}+\sigma)}{\lvert\nabla\rho_{1}\rvert}\right)_{+}\right]\frac{|\nabla\rho_{2}|^{2}}{\rho_{2}+\sigma}\,{\rm d}x\,{\rm d}t}_{\text{(II)}}.$
It is easy to see that the integrand in (II) is controlled by $|\nabla\sqrt{\rho}_{2}|^{2}\in L^{1}(\mathbb{R}^{d}\times(0,T))$. For (I) we use the Lipschitz condition (37) and we get
[…]
This shows that $\xi_{1}$ is uniformly bounded with respect to $\sigma$. Let us estimate $\xi_{2}$:
$\xi_{2}\leq$ $\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{2}}{\lvert\nabla\rho_{2}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\frac{\nabla\rho_{2}\nabla\rho_{1}}{\rho_{2}+\sigma}\,{\rm d}x\,{\rm d}t$
$-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{2}}{\lvert\nabla\rho_{2}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\frac{(\rho_{1}+\sigma)|\nabla\rho_{2}|^{2}}{(\rho_{2}+\sigma)^{2}}\,{\rm d}x\,{\rm d}t$
$\leq$ $\;\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{2}}{\lvert\nabla\rho_{2}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\frac{|\nabla\rho_{1}|^{2}}{\rho_{1}+\sigma}\,{\rm d}x\,{\rm d}t$
$-\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{2}}{\lvert\nabla\rho_{2}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\frac{(\rho_{1}+\sigma)|\nabla\rho_{2}|^{2}}{(\rho_{2}+\sigma)^{2}}\,{\rm d}x\,{\rm d}t$
$\leq$ $\;{\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{2}}{\lvert\nabla\rho_{2}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\frac{|\nabla\rho_{1}|^{2}}{\rho_{1}}\,{\rm d}x\,{\rm d}t}$
The integrand in the last expression is controlled by $|\nabla\sqrt{\rho_{1}}|\in L^{1}(\mathbb{R}^{d}\times(0,T))$. Therefore by dominated convergence we can take the limit $\sigma\rightarrow 0$ in (45) to obtain
$\int_{\mathbb{R}^{d}}$ $\left[\rho_{1}\log{\left(\frac{\rho_{1}}{\rho_{2}}\right)}-\rho_{1}+\rho_{2}\right](T)\,{\rm d}x{\leq\limsup_{\sigma\to 0}(\xi_{1}+\xi_{2})}\leq\;0.$
This implies that $\rho_{1}=\rho_{2}$.